### v1

Abstract
Biological intelligence uses a “multiscale competency architecture” (MCA). It exhibits adaptive, goal directed behaviour at all scales, from cells to organs to organisms. In contrast, machine intelligence is only adaptive and goal directed at a high level. Learned policies are passively interpreted using abstractions (e.g. arithmetic) embodied in static interpreters (e.g. x86). Biological intelligence excels at causal learning. Machine intelligence does not. Previous work showed causal learning follows from weak policy optimisation, which is hindered by presupposed abstractions in silico. Here we formalise MCAs as nested “agentic abstraction layers”, to understand how they might learn causes. We show that weak policy optimisation at low levels enables weak policy optimisation at high. This facilitates what we call “multiscale causal learning” and high level goal directed behaviour. We argue that by engineering human abstractions in silico we disconnect high level goal directed behaviour from the low level goal directed behaviour that gave rise to it. This inhibits causal learning, and we speculate this is one reason why human recall would be accompanied by feeling, and in silico recall not.
Keywords:
multiscale competency swarm intelligence consciousness.
1 Introduction
Biological intelligence is sample efficient. It is able to generalise or infer cause after only a few examples. Causal learning facilitates theory of mind, subjective experience and self awareness [bennett2023c, barron2016]. Our purpose here is to better understand how this capability might be engineered. Previous work established that causal learning requires one construct rather than presuppose variables and objects [bennett2023c]. By optimising for policies that express “weak” constraints on behaviour, relations between “contentless” [thompson2016] states are reduced to objects which cause affect [bennett2023b]. These claims were supported by proofs and experiments, but the resulting system is inefficient because it relies upon a high level, static abstraction layer. Intuitively, an abstraction layer is like a window through which only part of the world can be seen. The abstraction layer on which the aforementioned experiments were run conveys an inductive bias. It limits what sort of “causes” can efficiently be learned. How might we achieve efficient causal learning? What sort of abstraction layer would fix this problem?
1.0.1 Biological multiscale competency architecture:
Biology exhibits the capability we desire, so we look to biology for a more suitable abstraction layer. In the 70 years since psychoneural identity theory was proposed [sep-mind-identity], it has become evident that biological intelligence is not neatly confined to a neuronal abstraction layer [ciaunica2023]. Cells are now understood to be an agentic material [ball2023] and biology is slowly becoming a matter of engineering. It is now possible to create worms with two heads [morokuma2017], tadpoles with eyes on their backs that see [blackiston2013], and biological “xenobots” that perform simple tasks [levin2020]. Biology uses a “multiscale competency architecture” (MCA), in which lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [fields2020]. For example, cells have a “perceptual field”, and can co-operate with other cells in their perceptual field. This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [levin2024]. Cancer is an example of what happens when this fails [levin2021], as cells isolated from the informational structure of the collective lose their high level identity, and pursue lower level goals [davies2011]. MCA is relevant to AI because biological collectives exhibit high level goal directed behaviours far beyond the capability of any individual part [reid2011]. Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem. In other words, human intelligence is the high level goal directed behaviour of a collective of cells. If we want to engineer more human-like machine intelligence, then we should mimic an MCA.
1.0.2 Agentic Abstraction Layers:
High level goal directed behaviour can emerge from distributed systems like swarms of cells, but it is extremely difficult to precisely control this sort of emergent high level goal directed behaviour. Work by Ciaunica et. al. emphasised the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [ciaunica2023b]. Taking our cue from that approach, we propose formalising MCAs as nested and agentic abstraction layers. These abstraction layers are descriptions of behaviour, so we need not concern ourselves with the exact details of the system. It could be like biology, or something else. Our primary concern is how we might engineer and control efficient causal learning. A non-agentic, static abstraction layer passively interprets the abstraction layers above, like how a CPU based on the x86 instruction set architecture (ISA) would the behaviour of a python interpreter. By an agentic abstraction layer, we mean one that has its own agenda. It adapts, compelling the layers above to serve its own ends. Higher levels are compelled by lower to learn a policy “aligned” with theirs. Lower levels change their interpretation to enact [thompson2007] policy through higher levels, and are constrained by the resulting higher level behaviour. Information passes in both directions. Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate higher level goals.
1.0.3 Summary:
Do MCAs facilitate causal learning? How? Previous AGI research:
  1. 1.
    formalised enactive cognition, to describe goal directed behaviour in terms about which objective claims can be made [bennett2023b, bennett2023a],
  2. 2.
    proved that a weaker policies are optimal in terms of sample efficiency [^1] Also demonstrating a $110-500\%$ improvement in causal learning in experiments. [bennett2023b], causal learning, and
  3. 3.
    argued this facilitates theory of mind, reafference and self awareness [bennett2023c].
Here we:
  1. 1.
    examine the behaviour of MCAs in the context of this formalism, and show weaker policies at higher levels imply weaker policies at lower levels, facilitating what we call multiscale causal learning.
  2. 2.
    We argue a static ISA like x86 hinders causal learning in comparison, and
  3. 3.
    that lower levels “compel” higher. We speculate this may explain why human memory and feeling are intertwined. By disconnecting high level from low in silico, we make it possible to interpret without the impetus of “feelings”.
2 The Formalism
The following definitions are shared with [bennett2024b, bennett2023a]. More extensive explanations of these definitions are given there. We have endeavoured to make key aspects of this paper understandable without the formalism, so readers who wish to should simply skip to the next section. The definitions are here for those who wish to dig into the details.
Definition 1 (environment)
  • We assume a set $\Phi$ whose elements we call states.
  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all declarative programs (the powerset of $\Phi$).
  • By a truth or fact about a state $\phi$, we mean $f\in P$ such that $\phi\in f$.
  • By an aspect of a state $\phi$ we mean a set $l$ of facts about $\phi$ s.t. $\phi\in\bigcap l$. By an aspect of the environment we mean an aspect $l$ of any state, s.t. $\bigcap l\neq\emptyset$. We say an aspect of the environment is realised by state $\phi$ if it is an aspect of $\phi$.
Definition 2 (abstraction layer)
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite unless explicitly stated otherwise. If $\mathfrak{v}=P$, then we say that there is no abstraction.
  • $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$ is a set of aspects in $\mathfrak{v}$. We call $L_{\mathfrak{v}}$ a formal language, and $l\in L_{\mathfrak{v}}$ a statement.
  • We say a statement is true given a state iff it is an aspect realised by that state.
  • A completion of a statement $x$ is a statement $y$ which is a superset of $x$. If $y$ is true, then $x$ is true.
  • The extension of a statement $x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$. $E_{x}$ is the set of all completions of $x$.
  • The extension of a set of statements $X\subseteq{L_{\mathfrak{v}}}$ is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
  • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
(notation) $E$ with a subscript is the extension of the subscript222e.g. $E_{l}$ is the extension of $l$..
(intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer. The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table.
Definition 3 ($\mathfrak{v}$-task)
For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
  • ${I}_{\alpha}\subset L_{\mathfrak{v}}$ is a set whose elements we call inputs of $\alpha$.
  • ${O_{\alpha}}\subset E_{I_{\alpha}}$ is a set whose elements we call correct outputs of $\alpha$.
${I_{\alpha}}$ has the extension $E_{I_{\alpha}}$ we call outputs, and ${O_{\alpha}}$ are outputs deemed correct. $\Gamma_{\mathfrak{v}}$ is the set of all tasks given $\mathfrak{v}$.
[…]
(notation) If $\omega\in\Gamma_{\mathfrak{v}}$, then we will use subscript $\omega$ to signify parts of $\omega$, meaning one should assume $\omega=\langle{I}_{\omega},{O}_{\omega}\rangle$ even if that isn’t written.
(intuitive summary) To reiterate and summarise the above:
  • An input is a possibly incomplete description of a world.
  • An output is a completion of an input [def. 2].
  • A correct output is a correct completion of an input.
2.0.1 Learning and inference definitions:
Inference requires a policy and learning a policy requires a proxy. Strictly speaking, an organism is an embodied policy. However, organisms also adapt to particular tasks, so organisms also have and learn policies.
Definition 4 (inference)
  • A $\mathfrak{v}$-task policy is a statement $\pi\in L_{\mathfrak{v}}$. It constrains how we complete inputs.
  • $\pi$ is a correct policy iff the correct outputs $O_{\alpha}$ of $\alpha$ are exactly the completions $\pi^{\prime}$ of $\pi$ such that $\pi^{\prime}$ is also a completion of an input.
  • The set of all correct policies for a task $\alpha$ is denoted $\Pi_{\alpha}$. [^3] To repeat the above definition in set builder notation: ${\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$
Assume $\mathfrak{v}$-task $\omega$ and a policy $\pi\in L_{\mathfrak{v}}$. Inference proceeds as follows:
  1. 1.
    we are presented with an input ${i}\in{I}_{\omega}$, and
  2. 2.
    we must select an output $e\in E_{i}\cap E_{\pi}$.
  3. 3.
    If $e\in{O}_{\omega}$, then $e$ is correct and the task “complete”. $\pi\in{\Pi}_{\omega}$ implies $e\in{O}_{\omega}$, but $e\in{O}_{\omega}$ doesn’t imply $\pi\in{\Pi}_{\omega}$ (an incorrect policy can imply a correct output).
(intuitive summary) To reiterate and summarise the above:
  • A policy constrains how we complete inputs.
  • A correct policy is one that constrains us to correct outputs.
Definition 5 (learning)
  • A proxy $<$ is a binary relation on statements, and the set of all proxies is $Q$.
  • $<_{w}$ is the weakness proxy. For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$.
By the weakness of a statement, we mean the cardinality of its extension.
(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$ ($l$ is a cause of $\alpha$). We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
(probability of generalisation) We assume a uniform distribution over $\Gamma_{\mathfrak{v}}$. If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to $\alpha$ is less than the probability that $l_{2}$ generalizes to $\alpha$.
(sample efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more sample efficient than $<_{b}$ iff
$\left(\sum_{(l_{1},l_{2})\in\mathfrak{app}}\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{a}l_{2})\rvert-\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{b}l_{2})\rvert\right)<0$
(optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal. This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [bennett2023b]).
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child, to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is $<_{w}$.
Definition 6 (uninstantiated-tasks)
The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary). For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a highest level child $\omega\sqsubset\rho$ which is also a $\mathfrak{v}^{\prime}$-task. We call $\lambda_{\rho}$ an uninstantiated-task, and $\lambda_{1}\sqsubset\lambda_{2}$ iff $\lambda_{1}(P)\sqsubset\lambda_{2}(P)$.
Definition 7 (utility of intelligence)
Utility is the difference in weakness between the weakest and strongest correct policies of a task. The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
(intuitive summary) Uninstantiated tasks let us represent the same task in different abstraction layers. We can then use utility to measure how “good” an abstraction layer is for a particular task. This recognises that abstraction layers are not value-neutral. Each is better or worse for some tasks. Only the absence of abstraction is entirely neutral.
3 Arguments
We must first explain how $\mathfrak{v}$-tasks can represent MCAs. We must explain how different layers of the MCA are formed, and how “the behaviour of subunits percolates up” to higher level goal directed behaviours, which in turn “constrains and facilitates” lower level behaviours [levin2024]. Once that is done we show how an MCA can facilitate more efficient causal learning than a static ISA. We conceive of an MCA as a causal ‘chain’ where each node is the goal directed behaviour of the node that preceded it. For example, cells form organs which form organisms.
$cells$$organs$$organism$
We make no distinction between an object and its behaviour. Embodied and embedded, behaviour is limited to a particular vocabulary $\mathfrak{v}$. The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$. The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t. $\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$. Intuitively, this just means the two cells are parts of a larger whole. If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$). Note this identity is a precursor to the “causal identity for self” [bennett2023c] that differentiates observation from intervention in the manner of reafference [barron2016].
$\alpha_{1}$$\alpha_{2}$$\omega$
If the parts of a larger whole learn, then the probability that they learn a shared policy is maximised if they use the weakness proxy. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [bennett2023b, prop. 1,2]. For our cell example, weak policy at the level individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. In other words the construction of a weak policy is not just enabled by an MCA, but is required to form one at all. To explain how higher levels constrain lower, consider that higher levels in an MCA are just an abstraction; an emergent property we ascribe to the behaviour of the lowest level. For example, a collective of cells remains cells whether we call it an organ or not. A collective does not exist in the absence of its parts. $\omega$ and $\alpha_{1}$ and $\alpha_{2}$ all exist in the same vocabulary. If the outputs of $\alpha_{1}$ affect the inputs of $\alpha_{2}$ and vice versa, then the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The collective constrains behaviour at the lower level. We point this out to show how a $\mathfrak{v}$-task can represent an MCA. Just as the weakness proxy [bennett2023b, proposition 1,2] facilitates causal learning at a single scale [bennett2023c], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.
Intervening in the causal chain: We now illustrate how it is possible to “intervene” in the causal chain to disconnect high level behaviour from low level. We show this reduces “utility” (def. 7) for tasks at higher levels, and thus the efficiency of causal learning. That a collective of cells is an “organ” is just an abstract classification we ascribe to the collective’s behaviour. We’re interested in understanding how MCAs relate to abstraction layers in a computer. Just as an organ is really just cells, an “abstraction layer” is a property we ascribe to the emergent behaviour of many parts interacting. It appears to us as a second order system necessarily implied by the outputs $\mathfrak{v}$-tasks at different scales, because we see the whole rather than the parts. In the above example, $\omega$ is just $\alpha_{1}$ and $\alpha_{2}$ together, but to understand how we might “intervene” in the causal chain and disconnect emergent properties from the system which gave rise to them, we need to abstract away the individual cells and focus on the consequences of their behaviour. A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$. We use this fact to justify representing policies abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level. At each successive level, “correct” gets more specific. Because everything that happens in a higher level is correct according to lower levels, each higher level must further constrain “correct” if it is to exist in any meaningful sense.
$\lambda^{i-1}$$\lambda^{i}$$\lambda^{i+1}$
To construct layers we would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour which has emergent properties. We translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and we select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
Proposition 1 (propagation)
The greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t. $\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be (intuitively, this just means adaptability at higher levels implies adaptability at lower levels).
Proof
If $\mathfrak{a}\subset\mathfrak{b}$ then $\epsilon(\lambda^{i+1}(\mathfrak{a}))<\epsilon(\lambda^{i+1}(\mathfrak{b}))$, meaning if $\mathfrak{b}$ is the vocabulary at $i+1$, then it will be possible to construct weaker policies than if $\mathfrak{a}$ is the vocabulary (intuitively, a larger vocabulary enables a wider range of policies). We consider two policies $\pi^{i}_{\mathfrak{a}}$ and $\pi^{i}_{\mathfrak{b}}$ which could be the policy $\pi^{i}$ at $i$. If $\mathfrak{a}=\mathfrak{f}(E_{\pi^{i}_{\mathfrak{a}}})\subset\mathfrak{f}(E_{\pi^{i}_{\mathfrak{b}}})=\mathfrak{b}$, then $\pi^{i}_{\mathfrak{a}}<_{w}\pi^{i}_{\mathfrak{b}}$, meaning an enlarged vocabulary at $i+1$ implies a weaker policy at $i$. ∎
When we say we want efficient causal learning in computers, what we’re saying is that we want to mimic the adaptability of high level goal directed behaviour in humans. Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent, behavioural properties of that MCA.
$cells$$organs$$human$$abstractions$
When we create an abstraction layer like the ISA of a computer, we create a machine that mimics our more abstract behaviours. However, the ISA is fixed. It does not learn or adapt the way the lower levels of the human MCA does. It has no agency, the way cells do, because it is not adapting in pursuit of a goal. Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged. This interrupts the causal chain.
$human$$ISA$$abstractions$
As a result, there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels. Intuitively, if an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multiscale causal learning. We aren’t suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
Dropping Hume’s Guillotine: Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just abstractions we ascribe to the behaviour of the lowest level collective). Previous work argues learned behaviour must be motivated by attraction and repulsion from aspects of the environment [bennett2023c]. This eventually culminates in “feelings” at higher levels [bennett2023d]. Assuming that is true, then by intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that impetus. We might describe this as “dropping Hume’s Guillotine” to separate is from ought. We suggest that a computer is not a “syntactic engine”, as it is sometimes portrayed [kim2011]. Syntax has meaning to the computer, otherwise it would do nothing. However the “meaning” of an instruction is what it does, just as the “meaning” of a declarative program is its truth conditions. What a computer is, is meanings realised in circuitry by human hands. That is why it “interprets” syntax as it does. We speculate one reason computers may not “feel” and “experience” meaning as a human would is because there is a static, hard coded abstraction layer that allows high level goal directed behaviours to exist in the absence of the low level goal directed behaviours of which they are an emergent property. As a result, interpretation is fixed rather than fluid and motivated (e.g. by affect).
[…]

### v2

@@
-
+Multiscale Causal Learning
 Abstract
-Biological intelligence uses a “multiscale competency architecture” (MCA).
-It exhibits adaptive, goal directed behaviour at all scales, from cells to organs to organisms.
-In contrast, machine intelligence is only adaptive and goal directed at a high level.
-Learned policies are passively interpreted using abstractions (e.g.
-arithmetic) embodied in static interpreters (e.g.
-x86).
-Biological intelligence excels at causal learning.
-Machine intelligence does not.
-Previous work showed causal learning follows from weak policy optimisation, which is hindered by presupposed abstractions in silico.
-Here we formalise MCAs as nested “agentic abstraction layers”, to understand how they might learn causes.
-We show that weak policy optimisation at low levels enables weak policy optimisation at high.
-This facilitates what we call “multiscale causal learning” and high level goal directed behaviour.
-We argue that by engineering human abstractions in silico we disconnect high level goal directed behaviour from the low level goal directed behaviour that gave rise to it.
-This inhibits causal learning, and we speculate this is one reason why human recall would be accompanied by feeling, and in silico recall not.
+Biological intelligence is more sample-efficient than artificial intelligence (AI), learning from fewer examples.
+Here we answer why.
+Given data, there can be many policies which seem “correct” because they perfectly fit the data.
+However, only one correct policy could have actually caused the data.
+Sample-efficiency requires a means of discerning which.
+Previous work showed sample efficiency is maximised by weak-policy-optimisation (WPO); preferring policies that more weakly constrain what is considered to be correct, given finite resources.
+Biology’s sample-efficiency demonstrates it is better at WPO.
+To understand how, we formalise the “multiscale-competency-architecture” (MCA) observed in biological systems, as a sequence of nested “agentic-abstraction-layers”.
+We show that WPO at low levels enables synthesis of weaker policies at high.
+We call this “multiscale-causal-learning”, and argue this is how we might construct more scale-able, sample-efficient and reliable AI.
+Furthermore, a sufficiently weak policy at low levels is a precondition of collective policy at higher levels.
+The higher level “identity” of the collective is lost if lower levels use an insufficiently weak policy (e.g.
+cells may become isolated from the collective informational structure and revert to primitive behaviour).
+This has implications for biology, machine learning, AI-safety, and philosophy.
 keywords:
-multiscale competency swarm intelligence consciousness.
+multiscale competency, AI, causal learning, sample efficiency
 1 Introduction
-Biological intelligence is sample efficient.
-It is able to generalise or infer cause after only a few examples.
-Causal learning facilitates theory of mind, subjective experience and self awareness [bennett2023c, barron2016].
-Our purpose here is to better understand how this capability might be engineered.
-Previous work established that causal learning requires one construct rather than presuppose variables and objects [bennett2023c].
-By optimising for policies that express “weak” constraints on behaviour, relations between “contentless” [thompson2016] states are reduced to objects which cause affect [bennett2023b].
-These claims were supported by proofs and experiments, but the resulting system is inefficient because it relies upon a high level, static abstraction layer.
+Intelligence is often understood as predicting the outputs of a system, given inputs [1, 2, 3].
+In that frame, learning is the act of leveraging data [^1] A set of inputs and outputs known to correspond to a particular system.
+We also refer to this as training or historical data.
+to improve predictive accuracy.
+It is an optimisation problem.
+The cause of data is the policy [^2] By policy, we just mean a constraint on outputs and inputs.
+In other words, a policy is a declarative definition of behaviour.
+There is an equivalence between imperative and declarative programs [4] so intuitively, consider a policy to be a program that dictates behaviour.
+We use the word policy for the sake of consistency with related work and to avoid having to use two different terms for the same thing.
+We do not mean only the more narrow definition of policy used in reinforcement learning.
+which originally generated it.
+To borrow a philosophical term [5], it is the policy which was the original “causal intermediary” between outputs and inputs.
+In machine learning, training data are sets of examples of goal directed behaviour; an ostensive definition of how to complete a task.
+A learning algorithm synthesises a policy which “fits” training data.
+Intuitively, any policy that fits is “correct” for that training data.
+However, the fact that a policy is correct with respect to data is not enough information to determine whether that policy was the cause of that data.
+Data could be the result of any number of policies, and if one wishes to accurately predict outputs given inputs that aren’t in the training data, then one must identify that policy which originally generated the data.
+The true cause.
+The closer one’s policy is to the cause, the more accurately one will be able to predict outputs given new inputs one hasn’t seen before.
+Conversely, the better one is at identifying the likely causes of data, the less data one will require to learn how to complete the task.
+The less data one requires, the more “sample efficiently” one learns.
+1.1 Active Inference is Not Enough
+Active inference [6, 7] formalises how organisms learn and adapt.
+It describes living systems in terms of a machine learning approach called variational Bayes.
+Like other machine learning approaches, this treats learning as a process which synthesises a policy which happens fit data.
+A policy that is correct relative to historical data, but may not be the true cause of that data.
+Yet biological intelligence (BI) is far more sample efficient than artificial intelligence (AI).
+Organoids learn far quicker than deep reinforcement learning agents in silico [8].
+Yes, a robust AI will eventually learn cause and effect [9], but for now biology remains for more efficient.
+How is it that BI is so much more sample efficient than AI?
+This suggests there is more to the picture than active inference.
+BI is not just synthesising a policy which “fits” historical data; it is going a step further to figure out which among correct policies are likely to have caused historical data.
+How?
+1.2 Weak-Policy-Optimisation is Causal Learning
+Previous work already demonstrates how optimal sample efficiency [10] and causal learning [11] are achievable.
+Every policy implies a constraint on possible outputs and inputs; a constraint on how many possible tasks [^3] A task means a set of inputs and a corresponding set of outputs.
+A task is an ostensive definition of a goal or purpose.
+This is formally defined later in the paper.
+will be completed by this policy.
+A more specific policy completes fewer possible tasks, while a policy that implies a weaker constraint completes more possible tasks [12].
+Past research into causal learning and sample efficiency found that the correct policies most likely to have caused historical data imply the weakest [^4] Formal definitions for policy weakness are given in the next section.
+constraint [10, 11] [^5] See Bennett’s Razor and related proofs in [10] for sample efficiency, and [11] for causality arguments..
+Experiments demonstrated that a preference for weak policies correctly identifies the cause of historical data at $110-500\%$ the rate of a complexity based induction system [10].
+Gradient descent does not optimise for weak policies, which is one reason why deep learning is so sample inefficient [13].
+These results are further supported by the contravariance principle proposed by Cao and Yamins [14].
+They argue tasks that more tightly constrain the space of correct policies [^6] A larger, more varied set of inputs and outputs implies a smaller set of correct policies.
+make it easier to identify the policy that generated the inputs and outputs.
+This supports the claim that weaker policies are more likely to be the cause of behaviour, because a weaker policy is more likely to be a correct policy for tasks that more tightly constrain the set of correct policies.
+Given the sample efficiency exhibited by organoids [8], it would appear biology is far better at optimising for weak policies than current approaches to machine learning.
+Causal learning via weak-policy-optimisation [^7] Meaning learning using the weakness proxy as formally defined in the next section and in [3] (WPO) is clearly possible in silico, but existing implementations [10] are computationally expensive.
+One goal of our research is to better understand how the efficiency we observe in biology might be engineered.
+1.2.1 How Weak Policy Optimisation Constructs Causal Identities
+Previous work also established that causal learning requires one construct rather than presuppose variables and objects [11].
+Causal identities are learned by constructing $X$ such that $X$ causes $Y$, not by presupposing the environment is divided into objects including $X$ and then intervening to test if $X$ causes $Y$ [15].
+By optimising for policies that express “weak” constraints on behaviour, otherwise meaningless primitives888e.g.
+Declarative programs the truth value of which can be known, but for the truth conditions can never be known because environmental states are observable only partially and indirectly [3, 16, 15].
+or very low level abstractions may be clustered together to form “causal identities” for higher level abstractions, like variables [11].
+The “do” operator which is used to distinguish intervention from observation in causal inference problems can be substituted by introducing variables for interventions [17].
+That is why one can learn cause and effect via WPO [11], by constructing causal identities for interventions.
+1.2.2 How Causal Identities Relate to Abstraction Layers
+The construction of causal identities is a process by which an environment can be broken up into abstract classifications that are inherently meaningful (motivated and relevant to a goal).
+In philosophical terms, if one is motivated by attraction and repulsion from “contentless” [18] states [^9] Contentless meaning categorised, value-less, monolithic states with no discernible internal structure., then WPO is a means by which those states can be reduced to objects or events which cause the affect which motivates an organism [15].
+In computational terms, this amounts to constructing an abstraction layer that is most fit for a purpose [3, 16].
+For example, the x86 instruction set architecture (ISA) has hard coded interpretations of certain sorts, such as numbers and arithmetic operations [^10] These amount to policies by the formal definition we use [3, 16].
+As a result, the ISA can more efficiently address some problems than others.
+It is for this reason field programmable gate arrays exist as an alternative to static ISAs like x86; so that specific architectures can be constructed to more efficiently address specific problems.
+1.2.3 Static Abstraction Layers and Computational Dualism
+The ability to construct causal identities is easily hindered by the presupposition of static abstractions [11, 3].
+This is because one cannot construct weaker abstractions if one is forced to use a static set of high level abstractions, like numbers and arithmetic operations.
+This is symptomatic of problem called “computational dualism[3].
+Where 16th century Cartesian interactionist dualism proposed separate mental and physical substances interacting through the pineal gland, computational dualism is the concept of software distinct from hardware.
+For example, Hinton described software as “immortal”, existing infinitely many times across devices [19].
+According to the argument against computational dualism [3], immortal computations do not exist.
+There are only finitely many instances of mortal computations.
+Intuitively, when one loads software onto a computer, no object moves into the computer.
+All that happens is that the state of hardware changes.
+Addressing computational dualism revealed weakness to be the confounding variable that correlates simplicity with generalisation [16, Prop.
+2], allowing objective claims to be made regarding theoretical intelligent agents [3], addressing the subjectivity problems of complexity based induction systems in interactive settings [20].
+This is all relevant to us because static abstraction layers are based on computational dualism.
+They seem to suggest software is something other than the state of hardware.
+Something interchangeable.
+Move-able.
+This means static abstraction layers are necessarily one size fits all, rather than optimised for particular purpose.
+The inefficiency introduced at the abstraction layer affects everything built within it, including any software based machine learning systems.
 Intuitively, an abstraction layer is like a window through which only part of the world can be seen.
-The abstraction layer on which the aforementioned experiments were run conveys an inductive bias.
-It limits what sort of “causes” can efficiently be learned.
-How might we achieve efficient causal learning?
-What sort of abstraction layer would fix this problem?
-1.0.1 Biological multiscale competency architecture:
-Biology exhibits the capability we desire, so we look to biology for a more suitable abstraction layer.
-In the 70 years since psychoneural identity theory was proposed [sep-mind-identity], it has become evident that biological intelligence is not neatly confined to a neuronal abstraction layer [ciaunica2023].
-Cells are now understood to be an agentic material [ball2023] and biology is slowly becoming a matter of engineering.
-It is now possible to create worms with two heads [morokuma2017], tadpoles with eyes on their backs that see [blackiston2013], and biological “xenobots” that perform simple tasks [levin2020].
-Biology uses a “multiscale competency architecture” (MCA), in which lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [fields2020].
+It conveys an inductive bias.
+Presupposing the world is formed of numbers, objects and operations of a specific sort limits which cause and effect relations can efficiently be learned.
+The sample efficiency exhibited by biology suggests it uses a more versatile alternative to static abstraction layers based on computation dualism; one more suitable for WPO.
+1.3 Multiscale Competency Architecture
+Biology is slowly becoming a matter of engineering.
+It is now possible to create worms with two heads [21], tadpoles with eyes on their backs that see [22], and biological “xenobots” that perform simple tasks [23].
+As biology becomes more tractable, the similarities between biology and human artifice (e.g.
+computers) have become more obvious.
+Cells are now understood to be an agentic material [24] for the purpose of engineering.
+They are but one scale in the “multiscale competency architecture” (MCA) of biology.
+In an MCA lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [25].
+Adaptation takes place at every scale.
 For example, cells have a “perceptual field”, and can co-operate with other cells in their perceptual field.
-This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [levin2024].
-Cancer is an example of what happens when this fails [levin2021], as cells isolated from the informational structure of the collective lose their high level identity, and pursue lower level goals [davies2011].
-MCA is relevant to AI because biological collectives exhibit high level goal directed behaviours far beyond the capability of any individual part [reid2011].
+This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [26].
+Cancer happens when this fails [27].
+Cells isolated from the informational structure of the collective lose their high level identity, and pursue lower level goals [28].
+Biological collectives exhibit high level goal directed behaviours far beyond the capability of any individual part [29].
 Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem.
-In other words, human intelligence is the high level goal directed behaviour of a collective of cells.
-If we want to engineer more human-like machine intelligence, then we should mimic an MCA.
-1.0.2 Agentic Abstraction Layers:
+In other words, there is a great deal of evidence [26] to suggest that human intelligence is nothing more than the high level goal directed behaviour of a collective of cells.
+If we want sample-efficient AI, then we should mimic an MCA.
+2 Materials and Methods
+To summarise what has been covered so far:
+  1. 1.
+    Our goal is to understand why BI is so much more sample efficient (able not just to synthesise a policy that “fits” data, but to infer which among such correct seeming policies originally generated or caused the data).
+  2. 2.
+    As sample efficiency is maximised by WPO, our purpose is to understand how biology is able to synthesise weak policies so effectively.
+  3. 3.
+    A promising avenue appears to be the MCA [25] used by biological systems, instead of the static abstraction layers used in conventional computers (which exhibit the shortcomings of computational dualism [3]).
+Hypothesis: The agentic, nested structure of MCAs allows more effective and efficient WPO than would be possible within a static layer like x86.
+2.1 Agentic Abstraction Layers
+To begin to understand how this might be possible, we must formalise MCAs in rigorous, mathematical terms.
 High level goal directed behaviour can emerge from distributed systems like swarms of cells, but it is extremely difficult to precisely control this sort of emergent high level goal directed behaviour.
@@
 al.
-emphasised the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [ciaunica2023b].
+stressed the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [30].
 Taking our cue from that approach, we propose formalising MCAs as nested and agentic abstraction layers.
@@
 Our primary concern is how we might engineer and control efficient causal learning.
-A non-agentic, static abstraction layer passively interprets the abstraction layers above, like how a CPU based on the x86 instruction set architecture (ISA) would the behaviour of a python interpreter.
+A non-agentic, static abstraction layer passively interprets the abstraction layers above; like how a CPU based on the x86 instruction set architecture (ISA) interprets the behaviour of a python interpreter.
 By an agentic abstraction layer, we mean one that has its own agenda.
-It adapts, compelling the layers above to serve its own ends.
+It adapts.
+It compels the layers above to serve its own ends.
 Higher levels are compelled by lower to learn a policy “aligned” with theirs.
-Lower levels change their interpretation to enact [thompson2007] policy through higher levels, and are constrained by the resulting higher level behaviour.
+Lower levels change their interpretation to enact [31] policy through higher levels, and are constrained by the resulting higher level behaviour.
 Information passes in both directions.
-Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate higher level goals.
-1.0.3 Summary:
-Do MCAs facilitate causal learning? How? Previous AGI research:
-  1. 1.
-    formalised enactive cognition, to describe goal directed behaviour in terms about which objective claims can be made [bennett2023b, bennett2023a],
-  2. 2.
-    proved that a weaker policies are optimal in terms of sample efficiency [^1] Also demonstrating a $110-500\%$ improvement in causal learning in experiments. [bennett2023b], causal learning, and
-  3. 3.
-    argued this facilitates theory of mind, reafference and self awareness [bennett2023c].
-Here we:
-  1. 1.
-    examine the behaviour of MCAs in the context of this formalism, and show weaker policies at higher levels imply weaker policies at lower levels, facilitating what we call multiscale causal learning.
-  2. 2.
-    We argue a static ISA like x86 hinders causal learning in comparison, and
-  3. 3.
-    that lower levels “compel” higher.
-We speculate this may explain why human memory and feeling are intertwined.
-By disconnecting high level from low in silico, we make it possible to interpret without the impetus of “feelings”.
-2 The Formalism
-The following definitions are shared with [bennett2024b, bennett2023a].
-More extensive explanations of these definitions are given there.
-We have endeavoured to make key aspects of this paper understandable without the formalism, so readers who wish to should simply skip to the next section.
-The definitions are here for those who wish to dig into the details.
-Definition 1 (environment)
+Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate the pursuit of higher level goals.
+2.2 Definitions
+We begin with a formalism [3] of enactive cognition [31].
+We will use it to formalise MCAs in the next section.
+More extensive explanations of these definitions are given elsewhere [16, 3], and we have endeavoured to make key aspects of this paper understandable without the formalism; readers who wish to should simply skip to the next section.
+Definition 1. [environment].
   • We assume a set $\Phi$ whose elements we call states.
@@
 We say an aspect of the environment is realised by state $\phi$ if it is an aspect of $\phi$.
-Definition 2 (abstraction layer)
+Definition 2. [abstraction layer] .
   • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer.
-The vocabulary is finite unless explicitly stated otherwise.
+The vocabulary is finite unless stated otherwise [^11] This is because a bounded system can contain only a finite amount of information [32] in a spatially extended environment like our own [16]..
 If $\mathfrak{v}=P$, then we say that there is no abstraction.
@@
   • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
-(notation) $E$ with a subscript is the extension of the subscript222e.g. $E_{l}$ is the extension of $l$..
+(notation) $E$ with a subscript is the extension of the subscript121212e.g. $E_{l}$ is the extension of $l$..
 (intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer.
 The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table.
-Definition 3 ($\mathfrak{v}$-task)
-For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
+Definition 3. [$\mathfrak{v}$-task] For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
   • ${I}_{\alpha}\subset L_{\mathfrak{v}}$ is a set whose elements we call inputs of $\alpha$.
@@
   • An input is a possibly incomplete description of a world.
-  • An output is a completion of an input [def. 2].
+  • An output is a completion of an input [def. 2.2].
   • A correct output is a correct completion of an input.
@@
 However, organisms also adapt to particular tasks, so organisms also have and learn policies.
-Definition 4 (inference)
+Definition 4. [inference].
   • A $\mathfrak{v}$-task policy is a statement $\pi\in L_{\mathfrak{v}}$. It constrains how we complete inputs.
@@
   • A correct policy is one that constrains us to correct outputs.
-Definition 5 (learning)
+Definition 5. [learning] .
   • A proxy $<$ is a binary relation on statements, and the set of all proxies is $Q$.
@@
 (optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal.
-This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [bennett2023b]).
+This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [10]).
 (intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy.
@@
 The most sample efficient proxy is $<_{w}$.
-Definition 6 (uninstantiated-tasks)
-The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary).
+Definition 6.
+[uninstantiated-tasks] The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary).
 For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a highest level child $\omega\sqsubset\rho$ which is also a $\mathfrak{v}^{\prime}$-task.
 We call $\lambda_{\rho}$ an uninstantiated-task, and $\lambda_{1}\sqsubset\lambda_{2}$ iff $\lambda_{1}(P)\sqsubset\lambda_{2}(P)$.
-Definition 7 (utility of intelligence)
-Utility is the difference in weakness between the weakest and strongest correct policies of a task.
+Definition 7.
+[utility of intelligence] Utility is the difference in weakness between the weakest and strongest correct policies of a task.
 The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
@@
 Only the absence of abstraction is entirely neutral.
-3 Arguments
+3 Arguments and Results
 We must first explain how $\mathfrak{v}$-tasks can represent MCAs.
-We must explain how different layers of the MCA are formed, and how “the behaviour of subunits percolates up” to higher level goal directed behaviours, which in turn “constrains and facilitates” lower level behaviours [levin2024].
-Once that is done we show how an MCA can facilitate more efficient causal learning than a static ISA.
+We must explain how different layers of the MCA are formed; how “the behaviour of subunits percolates up” to determine higher level behaviour, which in turn “constrains and facilitates” lower level behaviours [26].
+Once that is done we show how an MCA can facilitate more efficient WPO than a static ISA.
 We conceive of an MCA as a causal ‘chain’ where each node is the goal directed behaviour of the node that preceded it.
@@
 We make no distinction between an object and its behaviour.
-Embodied and embedded, behaviour is limited to a particular vocabulary $\mathfrak{v}$.
+Embodied and embedded, behaviour is limited to a particular vocabulary $\mathfrak{v}$ (meaning there the number of possible configurations is finite).
 The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$.
@@
 If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$).
-Note this identity is a precursor to the “causal identity for self” [bennett2023c] that differentiates observation from intervention in the manner of reafference [barron2016].
+Note this identity is a precursor to the “causal identity for self” [11] that differentiates observation from intervention in the manner of reafference [33].
 $\alpha_{1}$$\alpha_{2}$$\omega$
-If the parts of a larger whole learn, then the probability that they learn a shared policy is maximised if they use the weakness proxy.
-This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [bennett2023b, prop.
+If the parts of a larger whole learn, then the probability that they learn a shared policy is maximised if they use the weakness proxy (what we have been referring to as WPO).
+This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [10, prop.
 1,2].
-For our cell example, weak policy at the level individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$.
-In other words the construction of a weak policy is not just enabled by an MCA, but is required to form one at all.
-To explain how higher levels constrain lower, consider that higher levels in an MCA are just an abstraction; an emergent property we ascribe to the behaviour of the lowest level.
+For our cell example, a sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$.
+In other words WPO is not just facilitated by an MCA, but required for high level behaviour.
+To explain how higher levels constrain lower, consider that higher levels in an MCA are just an abstraction; something we ascribe to the behaviour of the lowest level.
+For the sake of intuition; different levels are patterns within one larger pattern, and we simplify the larger pattern by dividing it up.
 For example, a collective of cells remains cells whether we call it an organ or not.
@@
 We point this out to show how a $\mathfrak{v}$-task can represent an MCA.
-Just as the weakness proxy [bennett2023b, proposition 1,2] facilitates causal learning at a single scale [bennett2023c], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.
+Just as the weakness proxy [10, proposition 1,2] facilitates causal learning at a single scale [11], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.
 Intervening in the causal chain: We now illustrate how it is possible to “intervene” in the causal chain to disconnect high level behaviour from low level.
 We show this reduces “utility” (def.
-7) for tasks at higher levels, and thus the efficiency of causal learning.
-That a collective of cells is an “organ” is just an abstract classification we ascribe to the collective’s behaviour.
+2.2.1) for tasks at higher levels, meaning reduced capacity for WPO.
 We’re interested in understanding how MCAs relate to abstraction layers in a computer.
-Just as an organ is really just cells, an “abstraction layer” is a property we ascribe to the emergent behaviour of many parts interacting.
-It appears to us as a second order system necessarily implied by the outputs $\mathfrak{v}$-tasks at different scales, because we see the whole rather than the parts.
-In the above example, $\omega$ is just $\alpha_{1}$ and $\alpha_{2}$ together, but to understand how we might “intervene” in the causal chain and disconnect emergent properties from the system which gave rise to them, we need to abstract away the individual cells and focus on the consequences of their behaviour.
+Just as an organ is really just cells, x86 is a classification we ascribe to the product of human behaviour.
+It is a higher level behaviour of an MCA that includes humans.
+To understand how we might “intervene” in the causal chain and disconnect higher level behaviours from the system which gave rise to them, we need to abstract away the lower levels.
+To relate this back to the cell example, we must abstract away individual cells and focus on the consequences of their behaviour.
 A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t.
-$\bigcap o=f$.
-We use this fact to justify representing policies abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$.
+$\bigcap o=f$ (we can substitute $o$ with a declarative program $f$ that has the same extension and thus truth conditions).
+We use this fact to justify representing policies as abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$.
 We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level.
@@
 If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$.
-$\pi^{i}$ implies behaviour which has emergent properties.
+$\pi^{i}$ implies behaviour.
 We translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$.
@@
 If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and we select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
-Proposition 1 (propagation)
+Theorem 1.
 The greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t.
 $\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be (intuitively, this just means adaptability at higher levels implies adaptability at lower levels).
-Proof
+Proof of Theorem 1.
 If $\mathfrak{a}\subset\mathfrak{b}$ then $\epsilon(\lambda^{i+1}(\mathfrak{a}))<\epsilon(\lambda^{i+1}(\mathfrak{b}))$, meaning if $\mathfrak{b}$ is the vocabulary at $i+1$, then it will be possible to construct weaker policies than if $\mathfrak{a}$ is the vocabulary (intuitively, a larger vocabulary enables a wider range of policies).
@@
 ∎
-When we say we want efficient causal learning in computers, what we’re saying is that we want to mimic the adaptability of high level goal directed behaviour in humans.
-Humans are not just part of MCAs.
-A human is an MCA.
-Human abstractions are emergent, behavioural properties of that MCA.
+Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent, behavioural properties of that MCA.
 $cells$$organs$$human$$abstractions$
@@
 It has no agency, the way cells do, because it is not adapting in pursuit of a goal.
+No graceful degradation.
 Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged.
@@
 After all, an increase in utility may increase the computational cost of searching for correct policies.
+Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [16], because bounded systems can contain only finite information [32].
 We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
-Dropping Hume’s Guillotine: Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just abstractions we ascribe to the behaviour of the lowest level collective).
-Previous work argues learned behaviour must be motivated by attraction and repulsion from aspects of the environment [bennett2023c].
-This eventually culminates in “feelings” at higher levels [bennett2023d].
-Assuming that is true, then by intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that impetus.
+3.1 Precondition for Collective Identity
+According to Cao and Yamins [14] tasks are more “difficult” when they have fewer correct policies, and “easier” when they have many correct policies.
+In our formalism, Cao and Yamins’ position implies $\mathfrak{v}$-task $\alpha$ is “easier” than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$.
+Collectives that exhibit high level goal directed behaviour that is “easier” have more correct high level policies to choose from.
+If the policy at a lower level $i-1$ changes such that the task at the higher level $i$ becomes “harder”141414i.e.
+by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in their being no correct policy at level $i$.
+If there are no longer any correct policies at level $i$ after a change in policy at level $i-1$, and if the system is robust and exhibits graceful degradation, then part of a collective may “break off” (loosening the constraint on correct policies, for both the breakaway part and the part left behind).
+In more formal terms, say we have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t.
+$\mathfrak{f}(E_{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$.
+Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t.
+$\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$.
+A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy.
+This is akin to “selective forgetting” [11] to cope with false negatives, but instead of forgetting data inconsistent with an otherwise good policy, we would “discard” a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that we obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t.
+$\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$.
+$\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of “correct” policies to choose from.
+This is analogous to cancer as cells isolated from a collective informational structure, that lose their high level identity, and pursue lower level goals [27, 28].
+If $\alpha$ is a cell, it can now revert to primitive transcriptional behaviours and act like cancer.
+4 Discussion
+We have formalised MCAs and shown the capacity for WPO at high levels implies WPO at low levels.
+Causal identities are constructed at different scales.
+Hence we call it multiscale-causal-learning (MCL).
+This seems to indicate MCA-like designs could result in more efficient and reliable AI in future.
+4.1 Dropping Hume’s Guillotine
+Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just classifications we ascribe to parts of the behaviour of the whole).
+Previous work argues learned behaviour must inevitably be motivated by attraction and repulsion from aspects of the environment [11].
+By intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that motivating impetus.
 We might describe this as “dropping Hume’s Guillotine” to separate is from ought.
-We suggest that a computer is not a “syntactic engine”, as it is sometimes portrayed [kim2011].
-Syntax has meaning to the computer, otherwise it would do nothing.
-However the “meaning” of an instruction is what it does, just as the “meaning” of a declarative program is its truth conditions.
-What a computer is, is meanings realised in circuitry by human hands.
-That is why it “interprets” syntax as it does.
-We speculate one reason computers may not “feel” and “experience” meaning as a human would is because there is a static, hard coded abstraction layer that allows high level goal directed behaviours to exist in the absence of the low level goal directed behaviours of which they are an emergent property.
-As a result, interpretation is fixed rather than fluid and motivated (e.g.
-by affect).
-[…]
+This relates to the hard problem of consciousness [34].
+According to our formalism, all abstractions constructed by living things are causal identities “goal directed” in the sense of being constructed in response to motivating affect.
+In dropping Hume’s Guilotine, high level human abstractions would become disconnected from the MCA that constructed them.
+It may be that, to a human or other MCA, all information amounts to statements of “ought” (motivated, value judgements) [15].
+Purely representational content (free of any goal directed motive or connotation) may exist only in the static abstraction layers we engineer.
+4.2 Future Research
+4.2.1 AI-Safety, Distributed and Cyberphysical-Systems
+The UK’s Advance Research and Invention Agency recently initiated a program called “Safeguarded AI” [35]; an ambitious research project that aims to erect a cyberphysical system meeting the standards laid out in Guaranteed Safe AI (GS) [36].
+This involves constructing advanced general AI of level ASL-4 [37] [^15] Meaning an artificial general intelligence [38] capable of autonomous research and development that can evaluate the outcomes of deploying less advanced AI, and safeguard humanity against “rogue” AI.
+in a cyberphysical system.
+That system seems intended to safeguard against threats from both within and without.
+The idea of a “rogue AI” from within this system is analogues to cancer, adopting an “unsafe” policy outside the scope of what is permitted by collective.
+This suggests we could formalise such a system as an MCA, in order to prove properties relevant to AI safety.
+Control of cyberphysical and distributed systems in general might be framed in terms of an MCA, in order to better understand how different levels in a hierarchy or different parts of a horizontal system constrain each-other.
+4.2.2 Active Inference
+The sample efficiency of organoids [8] reveals what active inference misses; WPO.
+Future research might integrate WPO with active inference.
+Surprise minimisation [7] could be considered an extremely weak policy in and of itself.
+However, what really matters is fitness [15], not surprise.
+In adopting machine learning as its foundation, active inference may have inadvertently dropped Hume’s Guillotine, emphasising the wrong goal.

### v3

@@
-Multiscale Causal Learning
+Are Biological Systems More Intelligent Than Artificial Intelligence?
 Abstract
-Biological intelligence is more sample-efficient than artificial intelligence (AI), learning from fewer examples.
-Here we answer why.
-Given data, there can be many policies which seem “correct” because they perfectly fit the data.
-However, only one correct policy could have actually caused the data.
-Sample-efficiency requires a means of discerning which.
-Previous work showed sample efficiency is maximised by weak-policy-optimisation (WPO); preferring policies that more weakly constrain what is considered to be correct, given finite resources.
-Biology’s sample-efficiency demonstrates it is better at WPO.
-To understand how, we formalise the “multiscale-competency-architecture” (MCA) observed in biological systems, as a sequence of nested “agentic-abstraction-layers”.
-We show that WPO at low levels enables synthesis of weaker policies at high.
-We call this “multiscale-causal-learning”, and argue this is how we might construct more scale-able, sample-efficient and reliable AI.
-Furthermore, a sufficiently weak policy at low levels is a precondition of collective policy at higher levels.
-The higher level “identity” of the collective is lost if lower levels use an insufficiently weak policy (e.g.
-cells may become isolated from the collective informational structure and revert to primitive behaviour).
-This has implications for biology, machine learning, AI-safety, and philosophy.
-keywords:
-multiscale competency, AI, causal learning, sample efficiency
+Is a biological self-organising system more ‘intelligent’ than an artificial intelligence?
+If so, why?
+We frame intelligence as adaptability, and explore this question using a mathematical formalism of enactive causal learning.
+We extend it to formalise the multilayer, multiscale, bottom-up distributed computational architecture of biological self-organisation.
+We then show that this architecture allows for more efficient adaptation than the static top-down interpreters typically used in computers.
+To put it provocatively, biology is more intelligent because cells adapt to provide a helpful inductive bias, and static interpreters do not.
+We call this multilayer-causal-learning.
+However it inherits a flaw of biological self-organisation.
+Cells become cancerous when isolated from the collective informational structure, reverting to primitive transcriptional behaviour.
+We show that, in the context of our formalism, failure states like cancer occur when systems are too tightly constrained by the abstraction layer in which they exist.
+This suggests control should be distributed (bottom-up rather than top-down) to ensure graceful degradation.
+We speculate about what this implies for systems in general, from machine learning hardware to human organisational and economic systems.
+Our result shows how we can design more robust systems and, though theoretical in nature, it lays a foundation for future empirical research.
 1 Introduction
-Intelligence is often understood as predicting the outputs of a system, given inputs [1, 2, 3].
-In that frame, learning is the act of leveraging data [^1] A set of inputs and outputs known to correspond to a particular system.
-We also refer to this as training or historical data.
-to improve predictive accuracy.
-It is an optimisation problem.
-The cause of data is the policy [^2] By policy, we just mean a constraint on outputs and inputs.
-In other words, a policy is a declarative definition of behaviour.
-There is an equivalence between imperative and declarative programs [4] so intuitively, consider a policy to be a program that dictates behaviour.
-We use the word policy for the sake of consistency with related work and to avoid having to use two different terms for the same thing.
-We do not mean only the more narrow definition of policy used in reinforcement learning.
-which originally generated it.
-To borrow a philosophical term [5], it is the policy which was the original “causal intermediary” between outputs and inputs.
-In machine learning, training data are sets of examples of goal directed behaviour; an ostensive definition of how to complete a task.
-A learning algorithm synthesises a policy which “fits” training data.
-Intuitively, any policy that fits is “correct” for that training data.
-However, the fact that a policy is correct with respect to data is not enough information to determine whether that policy was the cause of that data.
-Data could be the result of any number of policies, and if one wishes to accurately predict outputs given inputs that aren’t in the training data, then one must identify that policy which originally generated the data.
-The true cause.
-The closer one’s policy is to the cause, the more accurately one will be able to predict outputs given new inputs one hasn’t seen before.
-Conversely, the better one is at identifying the likely causes of data, the less data one will require to learn how to complete the task.
-The less data one requires, the more “sample efficiently” one learns.
-1.1 Active Inference is Not Enough
-Active inference [6, 7] formalises how organisms learn and adapt.
-It describes living systems in terms of a machine learning approach called variational Bayes.
-Like other machine learning approaches, this treats learning as a process which synthesises a policy which happens fit data.
-A policy that is correct relative to historical data, but may not be the true cause of that data.
-Yet biological intelligence (BI) is far more sample efficient than artificial intelligence (AI).
-Organoids learn far quicker than deep reinforcement learning agents in silico [8].
-Yes, a robust AI will eventually learn cause and effect [9], but for now biology remains for more efficient.
-How is it that BI is so much more sample efficient than AI?
-This suggests there is more to the picture than active inference.
-BI is not just synthesising a policy which “fits” historical data; it is going a step further to figure out which among correct policies are likely to have caused historical data.
-How?
-1.2 Weak-Policy-Optimisation is Causal Learning
-Previous work already demonstrates how optimal sample efficiency [10] and causal learning [11] are achievable.
-Every policy implies a constraint on possible outputs and inputs; a constraint on how many possible tasks [^3] A task means a set of inputs and a corresponding set of outputs.
-A task is an ostensive definition of a goal or purpose.
-This is formally defined later in the paper.
-will be completed by this policy.
-A more specific policy completes fewer possible tasks, while a policy that implies a weaker constraint completes more possible tasks [12].
-Past research into causal learning and sample efficiency found that the correct policies most likely to have caused historical data imply the weakest [^4] Formal definitions for policy weakness are given in the next section.
-constraint [10, 11] [^5] See Bennett’s Razor and related proofs in [10] for sample efficiency, and [11] for causality arguments..
-Experiments demonstrated that a preference for weak policies correctly identifies the cause of historical data at $110-500\%$ the rate of a complexity based induction system [10].
-Gradient descent does not optimise for weak policies, which is one reason why deep learning is so sample inefficient [13].
-These results are further supported by the contravariance principle proposed by Cao and Yamins [14].
-They argue tasks that more tightly constrain the space of correct policies [^6] A larger, more varied set of inputs and outputs implies a smaller set of correct policies.
-make it easier to identify the policy that generated the inputs and outputs.
+Biological systems self-organise, meaning their parts and the environment interact to produce spatiotemporal order [1, 2, 3, 4, 5, 6, 7].
+They interact with their environment to achieve homeostatic and reproductive goals.
+Some proactively learn [8], storing information about the past in order to anticipate future needs.
+To make accurate predictions in environments of sufficient complexity, a system must be able to learn cause and effect [9, 10, 11].
+For our purposes we consider the ‘intelligence’ of a system to be a measure of the resources it needs to correctly identify cause and effect [12, 13, 14, 15, 16] [^1] There are many definitions of intelligence, the merits of which are beyond the scope of this paper.
+Ours is primarily informed by [12], which is adapted from Legg and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’.
+The more efficient the learning system, the wider the range of environments in which it can satisfy goals.
+It is not possession of skills, but an ability to acquire skills with limited resources.
+Being an embodied formalism, this also accounts of energy and other resources consumed..
+It is a measure of how efficiently one adapts [^2] Primarily sample and energy efficiency.
+A ’sample’ is information used to learn.
+When a teacher provides an example of a math problem being solved, they are providing a sample.
+If student A requires 4 samples to understand a concept, and student B requires 5 samples, then A is more sample efficient than B..
+Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) of those that learn [17, 18, 19, 20].
+Contemporary approaches to machine learning are notoriously inefficient [21].
+They need vast quantities of data and energy learn tasks most humans can pick up after one or two examples.
+Even organoids, small collectives of cells grown in a lab, appear to be more ’intelligent’ in this sense than contemporary machine learning [22].
+Why is this the case?
+Are biological systems more intelligent, or is something else going on?
+Some suggest that the relative inefficiency of ML may be down to the fact that biological systems are active inference systems, and machine learning systems are typically not [23, 6, 24].
+Active inference is much more efficient [25, 26, 27], but it may only be part of the picture.
+As of 2024 active inference systems [^3] In particular the models submitted by Jack Cole and his employer, Minds AI.
+applied in the context of the ARC benchmark [14] [^4] For additional information, see https://arcprize.org for sample efficient adaptation still under-perform humans by a substantial margin [^5] Though it should also be noted that the Minds AI active inference system is one of the leading ARC submission for 2024, excluding OpenAI’s o3 which may or may not be an active inference system (the details have not been released at this time), and which is disqualified from ARC because it does not meet the computational efficiency requirements..
+Others hold that the reason biological systems appear to be more intelligent is because of an ‘inductive bias’ [14].
+This means the system embodies hard-coded foreknowledge of the environment, which predisposes it to efficient adaptation.
+If an inductive bias were the only reason for the discrepancy between biological self-organisation and machine learning, that would suggest biological systems are not more intelligent than artificial intelligence, but more knowledgeable.
+To reiterate, we seek to explain why biological self organisation seems to be more sample (and energy) efficient than contemporary machine learning.
+Is there something biological systems do that allows them to be more ’intelligent’ than artificial intelligence in its contemporary form?
+We suggest the answer lies in how biological self organisation distributes control at a different scales, and levels of abstraction.
+To do so we must first give some background on causal learning which informs our hypothesis.
+1.1 Causality
+Cause and effect are typically represented as an acyclic graph.
+causeeffect
+Complex environments often complicate the ability to identify causes.
+For example, confounded events may seem to cause one another when in fact they are merely correlated.
+This can lead one to draw erroneous conclusions about cause and effect.
+not the causeeffectconfounder
+By actively intervening in the environment a system can experiment to isolate causes.
+To do so, a system must be able to differentiate between events it passively observes and events it has caused [10, 9, 28].
+Then, it can construct a causal graph.
+There are two ways to do that.
+One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relation between them.
+This works well in theory for artificial intelligence [29], but requires foreknowledge of objects and properties between which causal relations can be established.
+The alternative is to assume the causal relations (the edges between nodes in the graph), and then learn the variables (the abstract concepts).
+This needs only attraction and repulsion from physical states [30], otherwise known as valence.
+We take this latter approach.
+We can frame causal learning in computational terms.
+Learning nodes in a causal graph is like figuring out which computer program was used to generate a sequence.
+Assume we have fed a set of inputs into a computer, and for each input the computer responded with a particular output.
+The computer is running a ’policy’ program which maps inputs to outputs [^6] We’ll call all such programs policies for the sake of this paper.
+This should not be confused with the more specific notion of a policy from reinforcement learning [19]..
+That policy is the cause of the input-output history we’ve observed so far.
+We want to know the policy that caused historical outputs so that we can use it to predict future outputs.
+The problem is that more than one policy exists which could have produced the history we’ve observed, and each of those policies produce different outputs when we give the computer inputs it hasn’t encountered before.
+From a naive Bayesian perspective, all of them are equally likely to have been the true underlying cause.
+If we wish to know which policy actually caused the outputs, we need way to narrow down the possibilities even further.
+There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [31].
+One way is to can compare their form, and choose programs which are simpler.
+However, in an interactive setting simplicity is an unreliable indicator of plausibility [32, 31].
+The alternative is to compare policies by their function, in terms of how they behave given inputs for which we do not have historical data.
+Adaptive systems frequently satisfy the same functional constraints by different means [33], so we can compare different sorts by the constraints they satisfy.
+Each policy implies a constraint on possible outputs given a set of possible inputs.
+The weaker that constraint is, the greater the variety of future behaviours it permits.
+Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker of the two [16].
+In theoretical terms the policy most likely to have caused behaviour is the weakest, and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [34].
+This has been used to establish local upper bounds on embodied ‘intelligence’ [12].
+ML systems do not typically optimise for either weaker constraints on function, or simplicity of form.
+This at least partially explains their relative sample inefficiency [35, 34].
+However, even when ML systems are modified to optimise for weakness, their capacity to do so limited by the “abstraction layer”.
+An abstraction layer is like a causal isolator [36] between a software “mind” and its environment.
+For example, the Python programming language is interpreted by the Python interpreter.
+In the end, Python gets translated into machine code, which hardware like an Apple M2 system on a chip then interprets to physical effect based on the ARM instruction set architecture.
+An abstraction layer is just physical embodiment by another name.
+Different abstraction layers permit different upper bounds on policy weakness, and thus different upper bounds on embodied ’intelligence’ [12].
+In other words, physical embodiment imposes severe limitations, and the behaviour of software is determined entirely by the hardware and software abstraction layers in which it runs.
+Intelligence is primarily a matter hardware, not software.
+In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [37, 38, 39].
+Put another way an embodied policy is a physical structure, like a tool.
+A tool can serve more than one purpose.
+For example, a hammer is both a weapon and a paperweight.
+A more specific tool completes fewer possible tasks, while a tool that implies a weaker constraint completes more possible tasks [21].
+By preferring weaker policies, one avoids unnecessary assumptions and so adapts as efficiently as possible, meaning the weakest policies complete a ‘wider variety of tasks’ than all others [^7] As an informal epistemological razor, this means “explanations should be no more specific than necessary”.
+Intuitively, this is a similar idea to the principle of maximum entropy [40, 41], but it accounts for the subjectivity of complexity and the teleological nature of language [31].
+For further intuition, a corollary to the idea that a weaker policy is more likely to be the cause of data is the inventors paradox, or equivalently the contravariance principle proposed by Cao and Yamins [42].
+They argue tasks that more tightly constrain the space of correct policies make better training sets.
+A larger, more varied set of inputs and outputs implies a smaller set of correct policies.
+That makes it easier to identify the policy that generated the inputs and outputs.
 This supports the claim that weaker policies are more likely to be the cause of behaviour, because a weaker policy is more likely to be a correct policy for tasks that more tightly constrain the set of correct policies.
-Given the sample efficiency exhibited by organoids [8], it would appear biology is far better at optimising for weak policies than current approaches to machine learning.
-Causal learning via weak-policy-optimisation [^7] Meaning learning using the weakness proxy as formally defined in the next section and in [3] (WPO) is clearly possible in silico, but existing implementations [10] are computationally expensive.
-One goal of our research is to better understand how the efficiency we observe in biology might be engineered.
-1.2.1 How Weak Policy Optimisation Constructs Causal Identities
-Previous work also established that causal learning requires one construct rather than presuppose variables and objects [11].
-Causal identities are learned by constructing $X$ such that $X$ causes $Y$, not by presupposing the environment is divided into objects including $X$ and then intervening to test if $X$ causes $Y$ [15].
-By optimising for policies that express “weak” constraints on behaviour, otherwise meaningless primitives888e.g.
-Declarative programs the truth value of which can be known, but for the truth conditions can never be known because environmental states are observable only partially and indirectly [3, 16, 15].
-or very low level abstractions may be clustered together to form “causal identities” for higher level abstractions, like variables [11].
-The “do” operator which is used to distinguish intervention from observation in causal inference problems can be substituted by introducing variables for interventions [17].
-That is why one can learn cause and effect via WPO [11], by constructing causal identities for interventions.
-1.2.2 How Causal Identities Relate to Abstraction Layers
-The construction of causal identities is a process by which an environment can be broken up into abstract classifications that are inherently meaningful (motivated and relevant to a goal).
-In philosophical terms, if one is motivated by attraction and repulsion from “contentless” [18] states [^9] Contentless meaning categorised, value-less, monolithic states with no discernible internal structure., then WPO is a means by which those states can be reduced to objects or events which cause the affect which motivates an organism [15].
-In computational terms, this amounts to constructing an abstraction layer that is most fit for a purpose [3, 16].
-For example, the x86 instruction set architecture (ISA) has hard coded interpretations of certain sorts, such as numbers and arithmetic operations [^10] These amount to policies by the formal definition we use [3, 16].
-As a result, the ISA can more efficiently address some problems than others.
-It is for this reason field programmable gate arrays exist as an alternative to static ISAs like x86; so that specific architectures can be constructed to more efficiently address specific problems.
-1.2.3 Static Abstraction Layers and Computational Dualism
-The ability to construct causal identities is easily hindered by the presupposition of static abstractions [11, 3].
-This is because one cannot construct weaker abstractions if one is forced to use a static set of high level abstractions, like numbers and arithmetic operations.
-This is symptomatic of problem called “computational dualism[3].
-Where 16th century Cartesian interactionist dualism proposed separate mental and physical substances interacting through the pineal gland, computational dualism is the concept of software distinct from hardware.
-For example, Hinton described software as “immortal”, existing infinitely many times across devices [19].
-According to the argument against computational dualism [3], immortal computations do not exist.
-There are only finitely many instances of mortal computations.
-Intuitively, when one loads software onto a computer, no object moves into the computer.
-All that happens is that the state of hardware changes.
-Addressing computational dualism revealed weakness to be the confounding variable that correlates simplicity with generalisation [16, Prop.
-2], allowing objective claims to be made regarding theoretical intelligent agents [3], addressing the subjectivity problems of complexity based induction systems in interactive settings [20].
-This is all relevant to us because static abstraction layers are based on computational dualism.
-They seem to suggest software is something other than the state of hardware.
-Something interchangeable.
-Move-able.
-This means static abstraction layers are necessarily one size fits all, rather than optimised for particular purpose.
-The inefficiency introduced at the abstraction layer affects everything built within it, including any software based machine learning systems.
-Intuitively, an abstraction layer is like a window through which only part of the world can be seen.
+See proofs 1 and 2 for more details..
+To complete a ‘wider variety of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses.
+That biological self-organising systems adapt efficiently suggests they are able to embody weaker policies than today’s ML systems.
+They attain higher local upper bounds on embodied intelligence [12].
+This suggests they have a means of optimising not just the choice of policy, but the embodied abstraction layer in which that policy exists.
+In a typical computer, the abstraction layers are static.
+For example, a Python interpreter does not learn and adapt to suit the program it is running.
+An abstraction layer is like a window through which only part of the world can be seen.
 It conveys an inductive bias.
-Presupposing the world is formed of numbers, objects and operations of a specific sort limits which cause and effect relations can efficiently be learned.
-The sample efficiency exhibited by biology suggests it uses a more versatile alternative to static abstraction layers based on computation dualism; one more suitable for WPO.
-1.3 Multiscale Competency Architecture
-Biology is slowly becoming a matter of engineering.
-It is now possible to create worms with two heads [21], tadpoles with eyes on their backs that see [22], and biological “xenobots” that perform simple tasks [23].
-As biology becomes more tractable, the similarities between biology and human artifice (e.g.
-computers) have become more obvious.
-Cells are now understood to be an agentic material [24] for the purpose of engineering.
-They are but one scale in the “multiscale competency architecture” (MCA) of biology.
+When we say that biological systems can optimise the abstraction layer, this is like saying they can “move” this window to learn a different inductive bias.
+The question is, what is the abstraction layer in the context of a biological self-organising system?
+We suggest the biological equivalent of abstraction layers are the layers of a “multiscale competency architecture” (MCA).
 In an MCA lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [43].
-Adaptation takes place at every scale.
-For example, cells have a “perceptual field”, and can co-operate with other cells in their perceptual field.
+Adaptation takes place at every scale, and cells are like an agentic material [44].
+Cells have a “perceptual field”, and can co-operate with other cells in their perceptual field.
 This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [45].
-Cancer happens when this fails [27].
-Cells isolated from the informational structure of the collective lose their high level identity, and pursue lower level goals [28].
-Biological collectives exhibit high level goal directed behaviours far beyond the capability of any individual part [29].
+Each cell pursues goals, and then ever larger groups of cells pursue collective goals at ever larger scales.
+Biological collectives exhibit high level goal directed behaviours beyond the capability of the individual parts [46].
 Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem.
-In other words, there is a great deal of evidence [26] to suggest that human intelligence is nothing more than the high level goal directed behaviour of a collective of cells.
-If we want sample-efficient AI, then we should mimic an MCA.
-2 Materials and Methods
-To summarise what has been covered so far:
-  1. 1.
-    Our goal is to understand why BI is so much more sample efficient (able not just to synthesise a policy that “fits” data, but to infer which among such correct seeming policies originally generated or caused the data).
-  2. 2.
-    As sample efficiency is maximised by WPO, our purpose is to understand how biology is able to synthesise weak policies so effectively.
-  3. 3.
-    A promising avenue appears to be the MCA [25] used by biological systems, instead of the static abstraction layers used in conventional computers (which exhibit the shortcomings of computational dualism [3]).
-Hypothesis: The agentic, nested structure of MCAs allows more effective and efficient WPO than would be possible within a static layer like x86.
-2.1 Agentic Abstraction Layers
-To begin to understand how this might be possible, we must formalise MCAs in rigorous, mathematical terms.
-High level goal directed behaviour can emerge from distributed systems like swarms of cells, but it is extremely difficult to precisely control this sort of emergent high level goal directed behaviour.
-Work by Ciaunica et.
-al.
-stressed the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [30].
-Taking our cue from that approach, we propose formalising MCAs as nested and agentic abstraction layers.
-These abstraction layers are descriptions of behaviour, so we need not concern ourselves with the exact details of the system.
-It could be like biology, or something else.
-Our primary concern is how we might engineer and control efficient causal learning.
-A non-agentic, static abstraction layer passively interprets the abstraction layers above; like how a CPU based on the x86 instruction set architecture (ISA) interprets the behaviour of a python interpreter.
+In computational terms this is a distributed system, like a von Neumann architecture [33].
+Our somewhat provocative claim is that biological self-organising systems are inherently more intelligent than artificial intelligence because their distributed, multilayer, multiscale architecture allows for higher local upper bounds on ’intelligence’.
+This is because it can shift abstraction layers to convey helpful inductive biases.
+Our results, though purely theoretical, lay a foundation for future empirical research.
+In section 2 introduce the formalism.
+In section 3 then argue this formalism represents an MCA and explain how.
+In the context of that formalism, we will formally show that an MCA allows for higher local upper bounds on embodied ‘intelligence’.
+We call this multilayer-causal-learning (MCL).
+In section 4 we explore how MCL may inherit the flaws of biological self-organising systems as well as their increased adaptability.
+We formalise a failure state analogous to cancer, and show that this state is more likely when a system is more tightly constrained.
+Finally, we relate these results to systems in general.
+We speculate about the possible implications, from more efficient hardware for machine learning to improved human organisational systems.
+2 Formalism and Proofs
+As part of our argument we attempt to describe MCAs in mathematically precise terms.
+Recent research has stressed the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [47].
+Taking our cue from that approach, we will formalise MCAs as a series of nested, adaptive, agentic abstraction layers.
+Enactivism frames cognition as an interaction between a self-organising system and its environment environment [48, 49, 50].
+We merely extend this process inwards, to formalise the interactions within the system itself.
+A non-agentic, static abstraction layer passively interprets the abstraction layers above.
 By an agentic abstraction layer, we mean one that has its own agenda.
-It adapts.
-It compels the layers above to serve its own ends.
-Higher levels are compelled by lower to learn a policy “aligned” with theirs.
+It adapts and compels the layers above to serve its own ends, while they each serve their own increasingly specific goals within the bounds of what lower levels permit.
+Higher levels are compelled by lower to learn a policy “aligned” with lower levels.
 Lower levels change their interpretation to enact [48] policy through higher levels, and are constrained by the resulting higher level behaviour.
@@
 Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate the pursuit of higher level goals.
-2.2 Definitions
-We begin with a formalism [3] of enactive cognition [31].
-We will use it to formalise MCAs in the next section.
-More extensive explanations of these definitions are given elsewhere [16, 3], and we have endeavoured to make key aspects of this paper understandable without the formalism; readers who wish to should simply skip to the next section.
-Definition 1. [environment].
+Except for the last definition, all the of the definitions that follow are from preceding work [12, 31].
+They formalise self-organisation in terms of constraints, like boundary conditions, as this lends itself to enactive depictions of intelligence [51].
+We have endeavoured to make key aspects of this paper understandable without the formalism.
+Readers who wish to should simply skip to the next section.
+{adjustwidth}
+11.6pt11.6pt
+Definition 1 (environment).
   • We assume a set $\Phi$ whose elements we call states.
-  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all declarative programs (the powerset of $\Phi$).
+  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all programs (the powerset of $\Phi$).
   • By a truth or fact about a state $\phi$, we mean $f\in P$ such that $\phi\in f$.
@@
 $\bigcap l\neq\emptyset$.
-We say an aspect of the environment is realised by state $\phi$ if it is an aspect of $\phi$.
-Definition 2. [abstraction layer] .
-  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer.
-The vocabulary is finite unless stated otherwise [^11] This is because a bounded system can contain only a finite amount of information [32] in a spatially extended environment like our own [16]..
-If $\mathfrak{v}=P$, then we say that there is no abstraction.
+We say an aspect of the environment is realised [^8] Realised meaning it is made real, or brought into existence.
+by state $\phi$ if it is an aspect of $\phi$.
+Definition 2 (abstraction layer).
+  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite.
   • $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$ is a set of aspects in $\mathfrak{v}$. We call $L_{\mathfrak{v}}$ a formal language, and $l\in L_{\mathfrak{v}}$ a statement.
@@
   • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
-(notation) $E$ with a subscript is the extension of the subscript121212e.g. $E_{l}$ is the extension of $l$..
-(intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer.
+(notation) $E$ with a subscript is the extension of the subscript999e.g. $E_{l}$ is the extension of $l$..
+$L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer.
 The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table.
-Definition 3. [$\mathfrak{v}$-task] For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
+Intuitively a sensorimotor system is an abstraction layer.
+Likewise, a computer (each statement asserting a state of the computer, or a part thereof).
+{adjustwidth}
+11.6pt11.6pt
+Definition 3 ($\mathfrak{v}$-task).
+For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
   • ${I}_{\alpha}\subset L_{\mathfrak{v}}$ is a set whose elements we call inputs of $\alpha$.
@@
 (notation) If $\omega\in\Gamma_{\mathfrak{v}}$, then we will use subscript $\omega$ to signify parts of $\omega$, meaning one should assume $\omega=\langle{I}_{\omega},{O}_{\omega}\rangle$ even if that isn’t written.
-(intuitive summary) To reiterate and summarise the above:
+To reiterate and summarise the above:
   • An input is a possibly incomplete description of a world.
-  • An output is a completion of an input [def. 2.2].
+  • An output is a completion of an input [def. 2].
   • A correct output is a correct completion of an input.
-2.2.1 Learning and inference definitions:
+2.1 Learning and inference definitions
 Inference requires a policy and learning a policy requires a proxy.
@@
 However, organisms also adapt to particular tasks, so organisms also have and learn policies.
-Definition 4. [inference].
+{adjustwidth}
+11.6pt11.6pt
+Definition 4 (inference).
   • A $\mathfrak{v}$-task policy is a statement $\pi\in L_{\mathfrak{v}}$. It constrains how we complete inputs.
@@
 $\pi\in{\Pi}_{\omega}$ implies $e\in{O}_{\omega}$, but $e\in{O}_{\omega}$ doesn’t imply $\pi\in{\Pi}_{\omega}$ (an incorrect policy can imply a correct output).
-(intuitive summary) To reiterate and summarise the above:
+To reiterate and summarise the above:
   • A policy constrains how we complete inputs.
   • A correct policy is one that constrains us to correct outputs.
-Definition 5. [learning] .
-  • A proxy $<$ is a binary relation on statements, and the set of all proxies is $Q$.
-  • $<_{w}$ is the weakness proxy. For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$.
-By the weakness of a statement, we mean the cardinality of its extension.
-(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$ ($l$ is a cause of $\alpha$).
+{adjustwidth}
+11.6pt11.6pt
+Definition 5 (learning).
+A proxy $<$ is a binary relation on statements.
+In this paper we use only one proxy, called the weakness proxy, which compares the cardinality of a statement’s extension.
+For statements $l_{1},l_{2}$ we have $l_{1}<l_{2}$ iff $\lvert Z_{l_{1}}\rvert<\lvert Z_{l_{2}}\rvert$.
+Whenever we use $<$ to compare statements, we are referring to the aforementioned weakness proxy.
+(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$.
 We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
@@
 If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to $\alpha$ is less than the probability that $l_{2}$ generalizes to $\alpha$.
-(sample efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more sample efficient than $<_{b}$ iff
+(efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more efficient than $<_{b}$ iff
 $\left(\sum_{(l_{1},l_{2})\in\mathfrak{app}}\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{a}l_{2})\rvert-\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{b}l_{2})\rvert\right)<0$
-(optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal.
-This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [10]).
-(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy.
+(optimal proxy) There is no proxy more efficient than weakness. The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [34]).
+Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy.
 Humans typically learn from “examples”.
 An example of a task is a correct output and input.
-A collection of examples is a child task, so “learning” is an attempt to generalise from a child, to one of its parents.
-The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy.
-The most sample efficient proxy is $<_{w}$.
-Definition 6.
-[uninstantiated-tasks] The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary).
+A collection of examples is a child task, so “learning” is an attempt to generalise from a child to one of its parents.
+The lower level the child from which an agent generalises to parent, the “faster” it learns (it chooses policies that complete a wider variety of tasks, and thus are more sample and energy efficient choices), the more efficient the proxy.
+The most efficient proxy is weakness (see proofs 1 and 2, or [34]), which is why we’re using it here.
+{adjustwidth}
+11.6pt11.6pt
+Definition 6 (uninstantiated-tasks).
+The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary).
 For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a highest level child $\omega\sqsubset\rho$ which is also a $\mathfrak{v}^{\prime}$-task.
 We call $\lambda_{\rho}$ an uninstantiated-task, and $\lambda_{1}\sqsubset\lambda_{2}$ iff $\lambda_{1}(P)\sqsubset\lambda_{2}(P)$.
-Definition 7.
-[utility of intelligence] Utility is the difference in weakness between the weakest and strongest correct policies of a task.
+Definition 7 (utility of intelligence).
+Utility is the difference in weakness between the weakest and strongest correct policies of a task.
 The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
-(intuitive summary) Uninstantiated tasks let us represent the same task in different abstraction layers.
-We can then use utility to measure how “good” an abstraction layer is for a particular task.
+Uninstantiated tasks let us represent the same task in different abstraction layers.
+We can then use utility to measure how “good” an abstraction layer is for a particular task [12].
 This recognises that abstraction layers are not value-neutral.
@@
 Only the absence of abstraction is entirely neutral.
-3 Arguments and Results
-We must first explain how $\mathfrak{v}$-tasks can represent MCAs.
-We must explain how different layers of the MCA are formed; how “the behaviour of subunits percolates up” to determine higher level behaviour, which in turn “constrains and facilitates” lower level behaviours [26].
-Once that is done we show how an MCA can facilitate more efficient WPO than a static ISA.
-We conceive of an MCA as a causal ‘chain’ where each node is the goal directed behaviour of the node that preceded it.
-For example, cells form organs which form organisms.
+2.2 Formalising Multilayer, Multiscale Architectures
+The following definition is explained at length over the remainder of this paper. Refer back here as needed.
+{adjustwidth}
+11.6pt11.6pt
+Definition 8 (multiscale competency architecture).
+  • The causal chain of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$.
+  • $\mathfrak{f}:2^{L_{\mathfrak{v}}}\rightarrow 2^{P}$ is a function that takes the extension $E_{\pi}$ of a policy $\pi\in L_{\mathfrak{v}}$, and returns a vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$.
+  • The state of the MCA is a sequence of policies $\langle\pi^{0},\pi^{1}...\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1}...\mathfrak{v}^{n}\rangle$ such that $\mathfrak{v}^{i+1}=\mathfrak{f}(E_{\pi^{i}})$ and $\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v^{i}})}$.
+In the absence of abstraction where the system is seen as nothing more than the sum of its parts, the MCA is just a task $\lambda^{0}(\mathfrak{v}^{0})$, allowing us to look at the system across scales of distribution.
+We say the MCA is over-constrained when there exists $i<n$ s.t.
+and $\Pi_{\lambda^{i}(\mathfrak{v^{i}})}=\emptyset$, and multilayer-causal-learning (MCL) occurs when the MCA is under-constrained and the proxy for learning is weakness.
+Note that the vocabulary is different at each level in the MCA, which means each has its own generational hierarchy of tasks.
+By a higher level of abstraction in the MCA, we mean a later task in the causal chain.
+2.3 Proofs
+Here we present three proofs.
+The first two demonstrate the optimality of the weakness proxy, and the last show that adaptability later in the causal chain necessitates adaptability earlier in the chain.
+Preliminary proof and experimental results showing weakness to be the optimal proxy were published in conference proceedings [34].
+However, the proofs contained minor errors.
+Most notably the earlier proof of sufficiency failed to account for task equivalency and so miscounted the total number of tasks, rather than counting the total number of equivalence classes [^11] Credit goes to Nora Belrose for spotting this error..
+This does not change the result, but for the sake of rigour these errors must be addressed.
+To that end, here we present more detailed alternatives to those proofs, which once again show the optimality of weakness as a proxy.
+Proposition 1 (sufficiency).
+Assume $\alpha\sqsubset\omega$.
+The weakness proxy sufficient to maximise the probability that a parent $\omega$ is learned from a child $\alpha$ [^12] Assume there exist correct policies for $\omega$, because otherwise there would be no point in trying to learn it..
+Proof.
+You’re given the definition of $\mathfrak{v}$-task $\alpha$ from which you infer a hypothesis $\pi\in{\Pi}_{\alpha}$. To learn $\omega$, you need $\pi\in\Pi_{\omega}$:
+  1. 1.
+    For every $\pi\in\Pi_{\alpha}$ there exists a $\mathfrak{v}$-task $\gamma_{\pi}\in\Gamma_{\mathfrak{v}}$ s.t.
+$O_{\gamma_{\pi}}=E_{\pi}$, meaning $\pi$ permits only correct outputs for that task regardless of input.
+We’ll call the highest level task $\gamma_{\pi}$ s.t.
+$O_{\gamma_{\pi}}=E_{\pi}$ the policy task of $\pi$.
+  2. 2.
+    $\omega$ is either the policy task of a policy in $\Pi_{\alpha}$, or a child thereof.
+  3. 3.
+    If a policy $\pi$ is correct for a parent of $\omega$, then it is also correct for $\omega$.
+Hence we should choose $\pi$ that has a policy task with the largest number of children.
+As tasks are uniformly distributed, that will maximise the probability that $\omega$ is $\gamma_{\pi}$ or a child thereof.
+  4. 4.
+    For the purpose of this proof, we say one task is equivalent [^13] This is because switching from $\beta$ to $\zeta$ s.t.
+$I_{\beta}\neq I_{\zeta}$ and $O_{\beta}=O_{\zeta}$ would be to pursue the same goal in different circumstances.
+This is because inputs are subsets of outputs, so both sets of inputs are implied by the outputs.
+$O_{\zeta}$ implies $I_{\beta}$ and $O_{\beta}$ implies $I_{\zeta}$ to another if it has the same correct outputs.
+  5. 5.
+    No two policies in $\Pi_{\alpha}$ have the same policy task [^14] Every policy task for policies of $\alpha$ is non-equivalent from the others..
+This is because all the policies in $\Pi_{\alpha}$ are derived from the same set inputs, $I_{\alpha}$.
+  6. 6.
+    The set of statements which might be outputs addressing inputs in ${I}_{\omega}$ and not ${I}_{\alpha}$, is $\overline{E_{{I}_{\alpha}}}=\{l\in L_{\mathfrak{v}}:l\notin E_{{I}_{\alpha}}\}$ [^15] This is because $E_{{I}_{\alpha}}$ contains every statement which is a correct output or an incorrect output, and $\overline{E_{{I}_{\alpha}}}$ contains every statement which could possibly be in $I_{\omega}$, $E_{I_{\omega}}$ and thus $O_{\omega}$..
+  7. 7.
+    For any given $\pi\in{\Pi}_{\alpha}$, the extension $E_{\pi}$ of $\pi$ is the set of outputs $\pi$ implies.
+The subset of $E_{\pi}$ which fall outside the scope of what is required for the known task $\alpha$ is $\overline{E_{{I}_{\alpha}}}\cap E_{\pi}$ [^16] This is because $E_{{I}_{\alpha}}$ is the set of all conceivable outputs by which one might attempt to complete $\alpha$, and so the set of all outputs that can’t be made when undertaking $\alpha$ is $\overline{E_{{I}_{\alpha}}}$ because those outputs occur given inputs that aren’t part of ${I}_{\alpha}$..
+  8. 8.
+    $\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|$ increases monotonically with $\left|E_{\pi}\right|$, because for all $e\in E_{\pi}$, $e\notin\overline{E_{{I}_{\alpha}}}$ iff $e\in E_{{I}_{\alpha}}$.
+This is because $L_{\mathfrak{v}}=\overline{E_{{I}_{\alpha}}}\cup E_{{I}_{\alpha}}$ and for all $\pi\in\Pi_{\alpha}$, $E_{\pi}\subseteq L_{\mathfrak{v}}$.
+  9. 9.
+    $2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|}$ is the number of non-equivalent parents of $\alpha$ to which $\pi$ generalises. It increases monotonically with the weakness of $\pi$.
+  10. 10.
+    Given $\mathfrak{v}$-tasks are uniformly distributed and $\Pi_{\alpha}\cap\Pi_{\omega}\neq\emptyset$, the probability that $\pi\in{\Pi}_{\alpha}$ generalises to $\omega$ is
+    $p(\pi\in{\Pi}_{\omega}\mid\pi\in{\Pi}_{\alpha},\alpha\sqsubset\omega)=\frac{2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|}}{2^{\left|\overline{E_{{I}_{\alpha}}}\right|}}$
+$p(\pi\in{\Pi}_{\omega}\mid\pi\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$ is maximised when $\left|E_{\pi}\right|$ is maximised.
+Recall from definition 4 that $<_{w}$ is the weakness proxy.
+For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$.
+$\pi$ that maximises $<_{w}$ will also maximise $p(\pi\in{\Pi}_{\omega}\mid\pi\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$.
+Hence the weakness proxy maximises the probability that [^17] Subsequently it also maximises the efficiency with which a parent $\omega$ is learned from a child $\alpha$.
+a parent $\omega$ is learned from a child $\alpha$.
+∎
+Proposition 2 (necessity).
+To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to use weakness as a proxy.
+Proof.
+Let $\alpha$ and $\omega$ be defined exactly as they were in the proof of prop. 1.
+  1. 1.
+    If $\pi\in{\Pi}_{\alpha}$ and $E_{{I}_{\omega}}\cap E_{\pi}={O}_{\omega}$, then it must be he case that ${O}_{\omega}\subseteq E_{\pi}$.
+  2. 2.
+    If $\left|E_{\pi}\right|<\left|{O}_{\omega}\right|$ then generalisation cannot occur, because that would mean that ${O}_{\omega}\not\subseteq E_{\pi}$.
+  3. 3.
+    Therefore generalisation is only possible if $\left|E_{\pi}\right|\geq\left|{O}_{\omega}\right|$, meaning a sufficiently weak hypothesis is necessary to generalise from child to parent.
+  4. 4.
+    For any two hypotheses $\pi_{1}$ and $\pi_{2}$, if $\mid E_{\pi_{1}}\mid<\mid E_{\pi_{2}}\mid$ then the probability $p(\left|E_{\pi_{1}}\right|\geq\left|{O}_{\omega}\right|)<p(\left|E_{\pi_{2}}\right|\geq\left|{O}_{\omega}\right|)$ because tasks are uniformly distributed.
+  5. 5.
+    Hence the probability that $\left|E_{{m}}\right|\geq\left|{O}_{\omega}\right|$ is maximised when $\left|E_{{m}}\right|$ is maximised.
+To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to select the weakest hypothesis.
+To select the weakest hypothesis, it is necessary to use the weakness proxy. ∎
+Proposition 3.
+The greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t. $\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be.
+Proof.
+If $\mathfrak{a}\subset\mathfrak{b}$ then $\epsilon(\lambda^{i+1}(\mathfrak{a}))<\epsilon(\lambda^{i+1}(\mathfrak{b}))$, meaning if $\mathfrak{b}$ is the vocabulary at $i+1$, then it will be possible to construct weaker policies than if $\mathfrak{a}$ is the vocabulary (intuitively, a larger vocabulary enables a wider range of policies).
+We consider two policies $\pi^{i}_{\mathfrak{a}}$ and $\pi^{i}_{\mathfrak{b}}$ which could be the policy $\pi^{i}$ at $i$.
+If $\mathfrak{a}=\mathfrak{f}(E_{\pi^{i}_{\mathfrak{a}}})\subset\mathfrak{f}(E_{\pi^{i}_{\mathfrak{b}}})=\mathfrak{b}$, then $\pi^{i}_{\mathfrak{a}}<_{w}\pi^{i}_{\mathfrak{b}}$, meaning an enlarged vocabulary at $i+1$ implies a weaker policy at $i$.
+∎
+3 Multiscale, Multilayer Causal Learning
+To begin, we must formally describe the architecture before we can consider its strengths and weaknesses in comparison to a typical computer.
+Hence we describe how $\mathfrak{v}$-tasks can describe the multiple scales and layers of abstraction in a biological system.
+We frame this first by looking at distribution of control within the system, and second by formalising the levels of abstraction at which we can view the system as a whole.
+Intuitively, this means we formalise different cross sections of the system.
+Looking at how control is distributed across different scales within the system we can see how the individual parts that make up a collective each possess their own policy or “identity”, and how that will (or will not) align with a shared collective policy or identity.
+This treats the system as nothing more than the sum of its parts.
+It is about self organisation, and how control is distributed within the resulting system.
+For example, individual neurons in the brain engage in predictive processing of eachother, and emulating this approach has yielded promising results in artificial neural networks [25, 26, 27].
+They form a network with a single coherent policy, and yet at the same time are individual neurons engaging in active inference.
+In contrast abstraction is about how the behaviour of cells enact an organ, and the behaviour of organs enacts an organism which enacts even more abstract behaviours like, for example, language.
+Each of these is a layer of abstraction.
+Layers exist in a causal ‘chain’ (a directed acyclic graph) where each layer (node) is the goal directed behaviour of the layer that preceded it.
+For example, cells form organs which form organisms [45].
 $cells$$organs$$organism$
-We make no distinction between an object and its behaviour.
-Embodied and embedded, behaviour is limited to a particular vocabulary $\mathfrak{v}$ (meaning there the number of possible configurations is finite).
+3.1 Distributing Self Organisation Across Scales
+The formalism we extend is enactive [12, 30], describing a self-organising system by describing what it does, rather than what it is.
+An embodied, embedded system can only do so many things, and so its behaviour is expressed using a finite vocabulary $\mathfrak{v}$ of declarative programs.
+These programs describe everything the system can do.
+For example if a system can be in a state of pain, then there are programs in $\mathfrak{v}$ which are “true” when the system is in such a state.
+Since states of the environment are mutually exclusive, $\mathfrak{v}$ implies an embodied formal language.
+By language here, we do not mean like the words humans use to communicate.
+We mean a language of concrete computations [52], where every statement in that language is rather than describes part of the state of a system (put another way, the interpreter is not assumed to be a UTM, but physical laws).
+For example, consider the language implied by a human being.
+If the human raises their arm, then they are expressing a statement in their embodied language.
+That statement then constrains what the body can do next, because everything that happens next begins from the point where the person had their arm raised [30].
+We frame self organisation in the same terms as the computer analogy we employed in the introduction.
+The behaviour of a self-organising system is goal directed, meaning there exists a constraint on its behaviour that defines that goal.
+The behaviour that satisfies the goal is expressed in the embodied formal language; it is a set of inputs and outputs we call a $\mathfrak{v}$-task.
+The inputs and outputs in the $\mathfrak{v}$-task are ‘fit’ or ‘correct’, the rest are not.
+Given these inputs, the system is fit if it embodies a statement that constrains behaviour to correct outputs only.
+We call statements that constrain behaviour policies.
+Policies constrain outputs given inputs, as in the computer analogy.
+Just as many different neural topologies can have the same functionality, many different policies can imply the same set of outputs given a particular set of inputs.
+A system learns if it can construct a policy from historical data which constrains behaviour to correct outputs (generalises beyond historical data), and once again the optimal choice of policy is the weakest.
+For brevity, when a system optimises for weak policies, we say it is engaged in weak policy optimisation (WPO).
+The above vocabulary can describe a distributed system.
+For example, an organ $\omega$ formed of two cells $\alpha_{1}$ and $\alpha_{2}$.
 The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$.
@@
 If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$).
-Note this identity is a precursor to the “causal identity for self” [11] that differentiates observation from intervention in the manner of reafference [33].
+Note this identity is a precursor to the “causal identity for self” [16] that, in its most basic form, differentiates observation from intervention in the manner of reafference [11, 30].
 $\alpha_{1}$$\alpha_{2}$$\omega$
-If the parts of a larger whole learn, then the probability that they learn a shared policy is maximised if they use the weakness proxy (what we have been referring to as WPO).
-This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [10, prop.
-1,2].
-For our cell example, a sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$.
-In other words WPO is not just facilitated by an MCA, but required for high level behaviour.
-To explain how higher levels constrain lower, consider that higher levels in an MCA are just an abstraction; something we ascribe to the behaviour of the lowest level.
-For the sake of intuition; different levels are patterns within one larger pattern, and we simplify the larger pattern by dividing it up.
-For example, a collective of cells remains cells whether we call it an organ or not.
-A collective does not exist in the absence of its parts.
-$\omega$ and $\alpha_{1}$ and $\alpha_{2}$ all exist in the same vocabulary.
-If the outputs of $\alpha_{1}$ affect the inputs of $\alpha_{2}$ and vice versa, then the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$.
-The collective constrains behaviour at the lower level.
-We point this out to show how a $\mathfrak{v}$-task can represent an MCA.
-Just as the weakness proxy [10, proposition 1,2] facilitates causal learning at a single scale [11], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.
-Intervening in the causal chain: We now illustrate how it is possible to “intervene” in the causal chain to disconnect high level behaviour from low level.
+A policy is like an identity described in the MCA literature [45].
+The probability that the parts of a system embody a collective policy is higher if they embody weaker policies.
+This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks (see proofs 1, 2 or [34]).
+Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$.
+A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$.
+A collective of cells remains cells regardless of whether we deem it to be an organ.
+The parts are still there.
+Formally the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary.
+Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$.
+The difference is a matter of scale.
+Just as the weakness proxy [34, proposition 1,2] facilitates causal learning at a single scale [16], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.
+However, this still just treats everything as a collective of cells.
+3.2 The Causal Chain of Abstraction
+From the enactive perspective, a human is something cells do, rather than cells are.
+To further clarify, we call a system complex when its behaviour is something we cannot easily comprehend or predict.
+However, just because we cannot comprehend and predict all of a system’s behaviour does not mean we cannot predict some of it.
+We suggest that so-called emergent properties are merely those aspects of a complex system’s behaviour which we can comprehend and predict.
+Aspects and patterns within patterns that remain comprehensible.
+For example, the overall pattern of a fractal like the Barnsley Fern [53] can be obvious, while the details elude comprehension.
+We now illustrate how behaviour forms a causal chain of layers of increasingly abstract behaviour, from cells to organs to organisms to the behaviour of those organisms (specifically human abstractions).
+We’ll then show how it is possible to ’intervene’ in that causal chain to disconnect more abstract behaviour from the less abstract behaviour that gave rise to it.
 We show this reduces “utility” (def.
-2.2.1) for tasks at higher levels, meaning reduced capacity for WPO.
-We’re interested in understanding how MCAs relate to abstraction layers in a computer.
-Just as an organ is really just cells, x86 is a classification we ascribe to the product of human behaviour.
+7) for tasks at higher levels, meaning reduced capacity for WPO.
+This is because the resulting distribution of control is altered when we intervene in the causal chain.
+Just as an organ boils down to a collection of cells, the collective of cells as a whole is an abstraction layer in which “organ” runs, like software.
+However, software is just a state of hardware [12].
+Put another way, “organ” is a classification we ascribe to the the behaviour of a cellular collective as a whole.
+As stated in the introduction, we’re interested in how MCAs relate to abstraction layers in a computer.
+x86 is a classification we ascribe to the product of human behaviour, a CPU.
 It is a higher level behaviour of an MCA that includes humans.
 To understand how we might “intervene” in the causal chain and disconnect higher level behaviours from the system which gave rise to them, we need to abstract away the lower levels.
-To relate this back to the cell example, we must abstract away individual cells and focus on the consequences of their behaviour.
+To relate this back to the cell example, we must abstract away individual cells and focus on the consequences of their collective behaviour, as a unified whole.
 A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t.
@@
 At each successive level, “correct” gets more specific.
-Because everything that happens in a higher level is correct according to lower levels, each higher level must further constrain “correct” if it is to exist in any meaningful sense.
+Because everything that happens in a higher level is correct according to lower levels, each higher level must more tightly constrain what is considered to be “correct” behaviour if that higher level is to exist in a meaningful sense.
 $\lambda^{i-1}$$\lambda^{i}$$\lambda^{i+1}$
@@
 If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and we select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
-Theorem 1.
-The greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t.
-$\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be (intuitively, this just means adaptability at higher levels implies adaptability at lower levels).
-Proof of Theorem 1.
-If $\mathfrak{a}\subset\mathfrak{b}$ then $\epsilon(\lambda^{i+1}(\mathfrak{a}))<\epsilon(\lambda^{i+1}(\mathfrak{b}))$, meaning if $\mathfrak{b}$ is the vocabulary at $i+1$, then it will be possible to construct weaker policies than if $\mathfrak{a}$ is the vocabulary (intuitively, a larger vocabulary enables a wider range of policies).
-We consider two policies $\pi^{i}_{\mathfrak{a}}$ and $\pi^{i}_{\mathfrak{b}}$ which could be the policy $\pi^{i}$ at $i$.
-If $\mathfrak{a}=\mathfrak{f}(E_{\pi^{i}_{\mathfrak{a}}})\subset\mathfrak{f}(E_{\pi^{i}_{\mathfrak{b}}})=\mathfrak{b}$, then $\pi^{i}_{\mathfrak{a}}<_{w}\pi^{i}_{\mathfrak{b}}$, meaning an enlarged vocabulary at $i+1$ implies a weaker policy at $i$.
-∎
+Theorem 3 shows that the greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t.
+$\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be.
+Intuitively, this just means adaptability later in the cause chain implies adaptability earlier in the chain.
 Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent, behavioural properties of that MCA.
-$cells$$organs$$human$$abstractions$
-When we create an abstraction layer like the ISA of a computer, we create a machine that mimics our more abstract behaviours.
-However, the ISA is fixed.
+$cells$$humans$$abstractions$
+When we create an abstraction layer like the ISA of a computer, we create a machine that embodies a narrow subset of our more abstract behaviours.
+On top of that ISA it can mimic all behaviours at higher levels of abstraction, because it is Turing complete.
+However, the ISA itself is fixed.
 It does not learn or adapt the way the lower levels of the human MCA does.
@@
 Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged.
-This interrupts the causal chain.
-$human$$ISA$$abstractions$
-As a result, there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels.
+This interrupts the causal chain and creates inefficiencies (which is why, for example, there is demand for application specific chipsets and field programmable gate arrays).
+$humans$$ISA$$abstractions$
+When we say the ISA is static, we mean there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels.
 Intuitively, if an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task.
@@
 When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA.
-In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multiscale causal learning.
+In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning.
+What proof 3 demonstrates is that adaptability at higher levels of abstraction demands adaptability at lower levels.
 We aren’t suggesting lower levels only optimise for utility at higher levels.
@@
 We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
-3.1 Precondition for Collective Identity
-According to Cao and Yamins [14] tasks are more “difficult” when they have fewer correct policies, and “easier” when they have many correct policies.
-In our formalism, Cao and Yamins’ position implies $\mathfrak{v}$-task $\alpha$ is “easier” than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$.
-Collectives that exhibit high level goal directed behaviour that is “easier” have more correct high level policies to choose from.
-If the policy at a lower level $i-1$ changes such that the task at the higher level $i$ becomes “harder”141414i.e.
+4 Drawbacks of Self Organisation
+We have proposed and explored some advantages of multilayer-causal-learning. Are there disadvantages? Do such systems also inherit the flaws of biological systems?
+Yes, they inherit a state analogous to cancer.
+In a biological system, when cells are isolated from the informational structure of their collective (i.e.
+an organ), they lose their high level identity.
+When that happens, they independently pursue their own lower level goals [55, 56].
+They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them.
+This is what is commonly known as cancer.
+In the context of our formalism, a state analogous to cancer occurs when a system is ’over-constrained’ by the abstraction layer in which it exists.
+To illustrate what we mean; the more difficult the task the more constrained the set of correct policies [42], meaning $\mathfrak{v}$-task $\alpha$ is easier than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$.
+Collectives that exhibit ‘easier’ high level goal directed behaviour in that sense have more correct high level policies to choose from.
+They are less tightly constrained.
+If the policy at a lower level $i-1$ changes such that the task at the higher level $i$ becomes “harder” in the sense of there existing fewer correct policies181818i.e.
 by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in their being no correct policy at level $i$.
-If there are no longer any correct policies at level $i$ after a change in policy at level $i-1$, and if the system is robust and exhibits graceful degradation, then part of a collective may “break off” (loosening the constraint on correct policies, for both the breakaway part and the part left behind).
+If there are no longer any correct policies at level $i$ after a change in policy at level $i-1$, then the only way the system can continue to function is for part of the collective to “splinter” or break away from the group, because that will loosen the constraint on correct policies for both the breakaway part and the collective it leaves behind.
 In more formal terms, say we have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t.
@@
 A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy.
-This is akin to “selective forgetting” [11] to cope with false negatives, but instead of forgetting data inconsistent with an otherwise good policy, we would “discard” a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that we obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t.
+This is akin to “selective forgetting” [16] of possible false negatives or positives in order to interpolate otherwise noisy data, but instead of forgetting data inconsistent with an otherwise good policy, we would discard a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that we obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t.
 $\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$.
 $\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of “correct” policies to choose from.
-This is analogous to cancer as cells isolated from a collective informational structure, that lose their high level identity, and pursue lower level goals [27, 28].
-If $\alpha$ is a cell, it can now revert to primitive transcriptional behaviours and act like cancer.
-4 Discussion
-We have formalised MCAs and shown the capacity for WPO at high levels implies WPO at low levels.
-Causal identities are constructed at different scales.
-Hence we call it multiscale-causal-learning (MCL).
-This seems to indicate MCA-like designs could result in more efficient and reliable AI in future.
-4.1 Dropping Hume’s Guillotine
+This is analogous to cancer.
+$\alpha$ is isolated and can now pursue goals which do not align with those of the collective.
+This suggests systems should be under-constrained to ensure graceful degradation or “sloppy fitness” [57], so it can maintain core functionality in as wide a variety of unanticipated circumstances as possible.
+An under-constrained system may be flexible enough to cope with unexpected inputs that more tightly restrict the set of correct policies.
+For example, if environment conditions are hostile or “difficult”, that would constitute an input that constrains the space of correct policies.
+The more difficult the environment, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters.
+Conversely, if control of a system is distributed among its parts at the lowest levels (i.e.
+a swarm rather than a central controller), then the destruction of its parts should not inhibit higher level function.
+There would be little reason for the parts of a system to splinter, as that would not greatly increase the number of available correct policies.
+5 Conclusions and Future Research
+We have formalised MCAs as a causal chain, and shown that the capacity for WPO later in the chain depends upon WPO earlier in the chain.
+This seems to indicate MCA-like designs could result in more adaptable and reliable AI in future.
+Indeed, the layered structure of artificial neural networks resembles exactly the sort of structure we have described, albeit constrained by the static abstraction layers in which an artificial neural network runs and without WPO.
+This may explain something of their relative success as a machine learning method.
+Future research might consider how hardware could be designed in a more distributed, multiscale fashion in order to facilitate MCL by fine tuning abstraction layers.
+In the context of the current hardware, adaptive abstraction layers might be considered, to facilitate greater efficiency.
+Beyond hardware, cyberphysical systems in general could be viewed through the lenses of abstraction and distribution we have described here, to ensure they are robust.
+We might consider how states analogous to cancer could arise in human organisational structures, or how power may be distributed throughout an organisation to ensure MCL.
+Free-market trade is a great example of how loose constraints have already facilitated adaptability in large collectives, and which is already being explored in the broader context of MCAs [58].
+Likewise, the delegation of greater control has been shown to improve the effectiveness of military organisations [59].
+Our claim boils down to a principle of design like an epistemological razor [34]; systems should be no more constrained than absolutely necessary.
+One concrete example of how this might be applied in future is the regulation of AI.
+The UK’s Advance Research and Invention Agency recently initiated a program called “Safeguarded AI” [60]; an ambitious research project that aims to erect a cyberphysical system meeting the standards laid out in Guaranteed Safe AI (GS) [61].
+This involves constructing advanced general AI of level ASL-4 [62] [^19] Meaning an artificial general intelligence [20] capable of autonomous research and development that can evaluate the outcomes of deploying less advanced AI, and safeguard humanity against “rogue” AI.
+in a cyberphysical system.
+That system seems intended to safeguard against threats from both within and without.
+The idea of a “rogue AI” from within this system is analogous to cancer, adopting an “unsafe” policy at odds with that of the collective.
+However if we view this system as MCA, then tight regulations may over-constrain the system trigger the cancerous state they are meant to prevent.
+This also might be applied to problems of a philosophical nature.
 Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just classifications we ascribe to parts of the behaviour of the whole).
@@
 This relates to the hard problem of consciousness [63].
-According to our formalism, all abstractions constructed by living things are causal identities “goal directed” in the sense of being constructed in response to motivating affect.
+All abstractions constructed by living things are “goal directed” in the sense that they are consequence of an evolutionary process [30].
 In dropping Hume’s Guilotine, high level human abstractions would become disconnected from the MCA that constructed them.
-It may be that, to a human or other MCA, all information amounts to statements of “ought” (motivated, value judgements) [15].
+It may be that, to a human or other MCA, all information amounts to statements of “ought” (motivated, value judgements).
 Purely representational content (free of any goal directed motive or connotation) may exist only in the static abstraction layers we engineer.
-4.2 Future Research
-4.2.1 AI-Safety, Distributed and Cyberphysical-Systems
-The UK’s Advance Research and Invention Agency recently initiated a program called “Safeguarded AI” [35]; an ambitious research project that aims to erect a cyberphysical system meeting the standards laid out in Guaranteed Safe AI (GS) [36].
-This involves constructing advanced general AI of level ASL-4 [37] [^15] Meaning an artificial general intelligence [38] capable of autonomous research and development that can evaluate the outcomes of deploying less advanced AI, and safeguard humanity against “rogue” AI.
-in a cyberphysical system.
-That system seems intended to safeguard against threats from both within and without.
-The idea of a “rogue AI” from within this system is analogues to cancer, adopting an “unsafe” policy outside the scope of what is permitted by collective.
-This suggests we could formalise such a system as an MCA, in order to prove properties relevant to AI safety.
-Control of cyberphysical and distributed systems in general might be framed in terms of an MCA, in order to better understand how different levels in a hierarchy or different parts of a horizontal system constrain each-other.
-4.2.2 Active Inference
-The sample efficiency of organoids [8] reveals what active inference misses; WPO.
-Future research might integrate WPO with active inference.
-Surprise minimisation [7] could be considered an extremely weak policy in and of itself.
-However, what really matters is fitness [15], not surprise.
-In adopting machine learning as its foundation, active inference may have inadvertently dropped Hume’s Guillotine, emphasising the wrong goal.
+In summary, we have formally described biological self organisation across distribution of scale and levels of abstraction, and shown adaptability at higher levels of abstraction depends upon adaptability at lower levels.
+We have described how a failure state analogous to cancer can occur when a system is too tightly constrained by the abstraction layer in which it exists (i.e.
+the task it faces is impossible).
+Though our result is of a purely theoretical nature, it illustrates how we may design more robust systems, and lays a foundation for future empirical research.

### v4

@@
 Abstract
-Is a biological self-organising system more ‘intelligent’ than an artificial intelligence?
+Are biological self-organising systems more ‘intelligent’ than artificial intelligence?
 If so, why?
-We frame intelligence as adaptability, and explore this question using a mathematical formalism of enactive causal learning.
-We extend it to formalise the multilayer, multiscale, bottom-up distributed computational architecture of biological self-organisation.
-We then show that this architecture allows for more efficient adaptation than the static top-down interpreters typically used in computers.
-To put it provocatively, biology is more intelligent because cells adapt to provide a helpful inductive bias, and static interpreters do not.
+We frame intelligence as adaptability, and explore this question using a mathematical formalism of causal learning.
+We compare systems by how they delegate control, illustrating how this applies with examples of computational, biological, human organisational and economic systems.
+We formally show the scale-free, dynamic, bottom-up architecture of biological self-organisation allows for more efficient adaptation than the static top-down architecture typical of computers, because adaptation can take place at lower levels of abstraction.
+Artificial intelligence rests on a static, human-engineered ‘stack’.
+It only adapts at high levels of abstraction.
+To put it provocatively, a static computational stack is like an inflexible bureaucracy.
+Biology is more ‘intelligent’ because it delegates adaptation down the stack.
 We call this multilayer-causal-learning.
-However it inherits a flaw of biological self-organisation.
+It inherits a flaw of biological systems.
 Cells become cancerous when isolated from the collective informational structure, reverting to primitive transcriptional behaviour.
-We show that, in the context of our formalism, failure states like cancer occur when systems are too tightly constrained by the abstraction layer in which they exist.
-This suggests control should be distributed (bottom-up rather than top-down) to ensure graceful degradation.
-We speculate about what this implies for systems in general, from machine learning hardware to human organisational and economic systems.
-Our result shows how we can design more robust systems and, though theoretical in nature, it lays a foundation for future empirical research.
+We show states analogous to cancer occur when collectives are too tightly constrained.
+To adapt to adverse conditions control should be delegated to the greatest extent, like the doctrine of mission-command.
+Our result shows how to design more robust systems and lays a mathematical foundation for future empirical research.
 1 Introduction
 Biological systems self-organise, meaning their parts and the environment interact to produce spatiotemporal order [1, 2, 3, 4, 5, 6, 7].
-They interact with their environment to achieve homeostatic and reproductive goals.
+They pursue homeostatic and reproductive goals.
 Some proactively learn [8], storing information about the past in order to anticipate future needs.
-To make accurate predictions in environments of sufficient complexity, a system must be able to learn cause and effect [9, 10, 11].
-For our purposes we consider the ‘intelligence’ of a system to be a measure of the resources it needs to correctly identify cause and effect [12, 13, 14, 15, 16] [^1] There are many definitions of intelligence, the merits of which are beyond the scope of this paper.
-Ours is primarily informed by [12], which is adapted from Legg and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’.
-The more efficient the learning system, the wider the range of environments in which it can satisfy goals.
-It is not possession of skills, but an ability to acquire skills with limited resources.
-Being an embodied formalism, this also accounts of energy and other resources consumed..
-It is a measure of how efficiently one adapts [^2] Primarily sample and energy efficiency.
-A ’sample’ is information used to learn.
+To make accurate predictions in complex environments, a system must learn cause and effect [9, 10, 11].
+We consider a system’s ‘intelligence’ to be a measure of how efficiently it does this [12, 13, 14, 15, 16] [^1] There are many definitions of intelligence, the merits of which are beyond the scope of this paper.
+Ours is primarily informed by [12], which frames intelligence as the ability to complete a wide range of tasks.
+This is an adapted from Legg and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’.
+Considering a ‘wide range of tasks’ accounts for embodiment, as well as environment and, being an embodied formalism, this also accounts of energy and other resources consumed.
+Intelligence by this definition is not possession of skills, but an ability to acquire skills with limited resources..
+Put another way, ‘intelligence’ is a measure of adaptability [^2] Primarily sample and energy efficiency.
+A ‘sample’ is information used to learn.
 When a teacher provides an example of a math problem being solved, they are providing a sample.
-If student A requires 4 samples to understand a concept, and student B requires 5 samples, then A is more sample efficient than B..
-Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) of those that learn [17, 18, 19, 20].
+If student A requires 4 samples to understand a concept, and student B requires 5 samples, then A is more sample efficient than B.
+Increased sample and energy efficiency means an increase in the total number of tasks one can complete, and thus adaptability..
+Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) is the study of machines that learn [17, 18, 19, 20].
 Contemporary approaches to machine learning are notoriously inefficient [21].
-They need vast quantities of data and energy learn tasks most humans can pick up after one or two examples.
-Even organoids, small collectives of cells grown in a lab, appear to be more ’intelligent’ in this sense than contemporary machine learning [22].
-Why is this the case?
-Are biological systems more intelligent, or is something else going on?
-Some suggest that the relative inefficiency of ML may be down to the fact that biological systems are active inference systems, and machine learning systems are typically not [23, 6, 24].
-Active inference is much more efficient [25, 26, 27], but it may only be part of the picture.
-As of 2024 active inference systems [^3] In particular the models submitted by Jack Cole and his employer, Minds AI.
-applied in the context of the ARC benchmark [14] [^4] For additional information, see https://arcprize.org for sample efficient adaptation still under-perform humans by a substantial margin [^5] Though it should also be noted that the Minds AI active inference system is one of the leading ARC submission for 2024, excluding OpenAI’s o3 which may or may not be an active inference system (the details have not been released at this time), and which is disqualified from ARC because it does not meet the computational efficiency requirements..
-Others hold that the reason biological systems appear to be more intelligent is because of an ‘inductive bias’ [14].
-This means the system embodies hard-coded foreknowledge of the environment, which predisposes it to efficient adaptation.
-If an inductive bias were the only reason for the discrepancy between biological self-organisation and machine learning, that would suggest biological systems are not more intelligent than artificial intelligence, but more knowledgeable.
-To reiterate, we seek to explain why biological self organisation seems to be more sample (and energy) efficient than contemporary machine learning.
-Is there something biological systems do that allows them to be more ’intelligent’ than artificial intelligence in its contemporary form?
-We suggest the answer lies in how biological self organisation distributes control at a different scales, and levels of abstraction.
-To do so we must first give some background on causal learning which informs our hypothesis.
-1.1 Causality
-Cause and effect are typically represented as an acyclic graph.
+They need vast quantities of data and energy to learn tasks a human can pick up after one or two examples.
+Even organoids [^3] By organoid we mean a collective of cells grown in lab, placed in a computational interface and life support system, and then ‘trained’ to communicate with conventional computing technology.
+appear more ‘intelligent’ in this sense [22].
+For example, an organoid outperformed conventional reinforcement learning systems at the game Pong, adapting with greater sample efficiency.
+That organoid was a highly tuned hybrid system whose behaviour was shaped by human-imposed boundary conditions.
+This is similar to how an acellular slime mold can be made to ‘solve’ shortest-path problems [23].
+These sorts of results are highly dependent upon the human imposed boundary conditions.
+Nevertheless, they raise an interesting question.
+Are biological self-organising systems more intelligent than artificial intelligence?
+If so, why?
+Some suggest that a key difference between biological systems and contemporary artificial intelligence is that the former are active inference systems, while the latter are typically not [24, 6, 25].
+It is true that active inference can improve efficiency [26, 27, 28].
+Active inference has been applied [^4] In particular the models submitted by Jack Cole and his employer, Minds AI.
+in the context of the ARC-AGI benchmark [14] [^5] For additional information, see https://arcprize.org.
+for sample efficient adaptation.
+Active inference systems perform comparatively well, but still under-perform humans [^6] Though it should also be noted that the Minds AI active inference system is one of the leading ARC submission for 2024, excluding OpenAI’s o3 which may or may not be an active inference system (the details have not been released at this time), and in any case o3 is extremely inefficient..
+Hence there may be more to the picture.
+Active inference aside, some hold biological systems are not more intelligent.
+They only seem to be because of a ‘strong inductive bias’ [14].
+This means the system embodies hard-coded foreknowledge of the environment, which predisposes it to adapt.
+To whatever extent this is true, it merely substitutes the question ‘what is intelligence?’ for ‘what is the inductive bias?’.
+To reiterate, we want to know why biological self-organisation seem more adaptive [^7] Sample and energy efficient.
+than contemporary AI.
+Is there something biological systems do that allows them to be more ‘intelligent’?
+Could there be more to the picture than active inference and an inductive bias?
+We suggest there is.
+A notable feature of biological self-organisation is the multiscale competency architecture (MCA).
+Adaptation takes place at every scale, from cells to organs to organisms to entire ecosystems [29].
+It is ‘scale-free’.
+Similar dynamics tend to repeat at different scales.
+Control is delegated.
+Cells are a “soft” material that has agency [30, 31].
+A cell has a perceptual field.
+A cell can co-operate with other cells in this perceptual field.
+This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [32].
+Each cell has its own goals and is intelligent in the sense of being able to reach the same goal by different means [33, 34].
+However each cell can also be part a collective of cells that in turn pursues its own collective goals.
+Smaller collectives can form larger collectives, and the pattern repeats at ever larger scales.
+Collectives are nested [35, 36].
+Each “thinks” in the enactive sense, interacting with its environment [37, 38, 39].
+Biological collectives can exhibit high level goal directed behaviour far beyond the capability of the parts of which they’re formed [40].
+Figure 1: The Barnsley Fern is an iterated function system fractal that exhibits self similarity [41].
+Notice how the same pattern repeats at different scales.
+This gives some visual intuition as to what ‘scale-free’ suggests, as we can see the same dynamics play out again and again at different scales.
+Our somewhat provocative claim is that biological self-organising systems are inherently more intelligent than contemporary implementations of artificial intelligence.
+We will argue this is because biological self-organisation delegates control to the lowest possible level of abstraction.
+We also present three proofs.
+The first two show that adaptability depends on the ability to embody “weak” constraints.
+The third proves the ability to embody weak constraints at higher levels of abstraction depends on the ability to embody weak constraints at lower levels.
+Our results, though purely theoretical, lay a foundation for future empirical research.
+2 Abstraction and Delegation of Systems
+Here we will briefly explain what we mean by abstraction and delegation.
+These are dimensions along which we may plot any system.
+For intuition, we will give examples of limit cases for computational, biological, human organisational and economic systems.
+2.1 Abstraction
+Abstraction is about how a system produces emergent, increasingly abstract behaviours.
+Highly abstract behaviours are a consequence of simpler goal directed behaviour at lower levels of abstraction.
+For example a collective of cells can enact an organ, and a system of organs can enact an organism [32] which enacts even more abstract behaviours like language or tools.
+Each of these is an ‘abstraction layer’ [12].
+Abstraction layers exist in a stack (which we represent as graph).
+Intuitively, this is a ‘chain’ of cause and effect where each link in the chain (an abstraction layer) is the goal directed behaviour of the link (layer) that preceded it.
+$cells$$organs$$organism$
+When we say we examine a system at a lower level of abstraction, we mean we look at a node lower in the stack (or earlier in the chain).
+In the case of biological systems lower levels of abstraction are formed of smaller or simpler parts [42], so looking at a lower level of abstraction is like ‘zooming in’ to look at a smaller scale.
+Conversely, rather than looking at low levels of abstraction like cells we can shift the level of abstraction to higher in the stack, and look at individual humans and how they form collectives.
+$soldiers$$squads$$platoon$
+A conventional computing system is a stack of abstraction layers. For example a predictive model typically relies on a machine learning library, which runs on CUDA, which runs on an Nvidia GPU.
+$GPU$$CUDA$$library$$model$
+2.2 Delegated Control
+Unlike biological systems a lower level of abstraction in a computer is not necessarily made up of smaller parts that form a collective.
+Looking at a lower level of abstraction is not like ‘zooming in’.
+For example, Python code is run by an interpreter, which can run on a computer.
+That interpreter is not a collective.
+It is not made up of parts capable of functioning independently.
+Delegating work is not the same thing as delegating control.
+For example, work can be delegated to GPUs, but if we remove a GPU from the computer it will not then pursue goals independently.
+It will sit, inert until once again integrated into a larger system.
+It requires no incentives.
+Control is imposed top-down, from higher to lower levels of abstraction.
+Conversely in a biological system each cell is independently capable of goal directed behaviour [31], assuming it remains alive.
+Control in a complex biological system begins bottom-up, before top-down.
+Control then propagates in both directions.
+Each and every part of a collective acts on its own, and those actions imply a ‘policy’.
+Each individual policy determines how that part interacts with the other parts, or more to the point their policies.
+These interactions cause the system as a whole to act, which implies a collective policy.
+If collective and individual policies align in service of homeostasis, then the system can persist.
+It has a collective identity or goal.
+That identity can constrain parts of the collective.
+For example, we might say the level of abstraction is cells, and then describe an organ as a collective policy of those cells.
+Likewise, we might look at the level of organs, and describe a human as a collective policy of those organs.
+Or we could describe the human as the collective policy of cells, and see the organs as a sort of middle management layer.
+Figure 2: Illustration of a decentralised system (left) vs a centralised system (right).
+Notice that if a central node from the system on the right is deleted, then large portions of the graph become disconnected.
+If a node is deleted from the left graph, then there would still be a path between all remaining nodes.
+The delegation of control necessitates decentralisation.
+A good example to understand the limit cases for delegation is an economy.
+In an ideal centrally planned economy control is concentrated at the highest level of abstraction.
+We can think of the people in the economy as a lower level of abstraction.
+Those people can adapt, but are not really allowed to because all decisions are imposed top-down.
+On the other extreme control is delegated to parts at the lowest level of abstraction, and conditions are such that control remains decentralised (i.e.
+there is some method of redistribution such as war or technological innovation).
+When we choose a level of abstraction, we choose a smallest unit of delegation.
+For example if the level of abstraction is humans, then we don’t worry about how control is delegated within a human.
+We just concern ourselves with what systems can be formed of humans.
+Like an economy.
+Of course, there are limits on what sort of physical forms can store or process information in service of a goal [43].
+This means there are limits to how much control can be delegated.
+Figure 3: Examples of limit cases for delegation at different levels of abstraction.
+2.3 Reasons to Suspect Delegation Facilitates Adaptation
+We will now give a brief explanation of why we believe delegation of control is relevant to adaptation.
+All else being equal the lower the level of abstraction at which the system can be adapted, the more it can be specialised.
+For example a program written in C can be far more efficient than a program written in Python [44].
+This is because the Python program has to be interpreted by the Python interpreter, which is a program written in C.
+By programming in C we simply cut out an abstraction layer.
+This gives the programmer more leeway to optimise the system, or make mistakes.
+Likewise another way to make computers more efficient is to use a field programmable gate array (FPGA) to write a program at the hardware level, instead of compiling that program into machine code and running it on a one-size-fits-all instruction set architecture like x86 [45].
+With each successive abstraction layer introduced, one is further removed from the ground truth to which one must adapt.
+For example software is nothing more than a state of hardware [12].
+It is a label we ascribe to the behaviour of hardware.
+It is a useful simplification that lets us consider what we want hardware to do in the abstract, but working at such a high level of abstraction can cause problems.
+Proposals for software ‘intelligence’ have been undermined by the fact that any claims regarding the behaviour of software are contingent upon hardware.
+The software doesn’t interact with the environment, the hardware does.
+This is because the ability to adapt is related to perceived complexity, which is entirely determined by the abstraction layer through which an agent interacts with its environment [46].
+Consequently an ‘objective’ measure of artificial intelligence [47, 12] must be ‘enactive’ [37] in that it hinges not just software but hardware and the local environment as well.
+To understand the difference between biological systems and artificial intelligence, we must account for the level of abstraction at which an adaptive process takes place.
+In distributed computational systems, like a drone swarm or network, delegating control can facilitate what is called ‘graceful degradation’ [48].
+For example, to control a drone swarm we could single out one drone to be the ‘leader’, and have it run a program that co-ordinates the other drones.
+However that would make the swarm vulnerable.
+If the leader drone were damaged, the swarm would de-cohere.
+It is a single point of failure.
+In contrast if we decentralise control, delegating to the level of individual drones, the swarm will remain coherent regardless of how many drones are disabled.
+In other words, the swarm ‘adapts’ to adverse conditions.
+The advantages of delegating control are found in human organisations as well.
+For example, in a military context there is a concept called ‘mission command’ [49].
+A culture of mission command empowers units to adapt the situation at hand, by granting them leeway to bend rules and orders in service of the overall mission.
+It loosens the constraints under which units must operate, delegating control.
+Officers at every level are trained to take the initiative and act independently.
+A well documented example of the effectiveness of mission command in facilitating adaptation is that of Nordbat 2, a joint peacekeeping force deployed to Bosnia in the early 90s.
+Mission command allowed individual units to deviate from convention, adapt to local conditions and adopt a more ‘trigger-happy’ disposition that proved to be substantially more effective than the more rigid, rule based approach taken by the Dutch forces operating in the same region [50].
+Each layer of management is like an abstraction layer.
+When control is delegated, feedback loops are shortened because layers of abstraction are removed between the controlling entity and the situation at hand.
+Greater degrees of adaptation are permitted at lower levels of abstraction.
+This all lends credence to the idea that delegation of control is an important feature of biological systems which is absent from mainstream contemporary artificial intelligence.
+3 What is Adaptation?
+We will formally establish that delegation of control makes a system more adaptable, all else being equal.
+Biological systems delegate adaptation down the stack, while computers typically do not.
+Hence we will argue biological systems are more intelligent, but first we must describe biological systems and computers in comparable terms.
+We will begin by discussing what it means to learn and adapt, in terms of cause and effect.
 causeeffect
-Complex environments often complicate the ability to identify causes.
-For example, confounded events may seem to cause one another when in fact they are merely correlated.
+To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [10, 9, 51].
+It is often difficult to determine cause.
+‘Confounding’ occurs when an event appears be the cause of the effect we’re trying to predict, but in truth is merely a correlated effect.
 This can lead one to draw erroneous conclusions about cause and effect.
-not the causeeffectconfounder
-By actively intervening in the environment a system can experiment to isolate causes.
-To do so, a system must be able to differentiate between events it passively observes and events it has caused [10, 9, 28].
+not the causeeffectcause
+By actively intervening in the environment a system can experiment to establish cause and effect.
 Then, it can construct a causal graph.
 There are two ways to do that.
-One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relation between them.
-This works well in theory for artificial intelligence [29], but requires foreknowledge of objects and properties between which causal relations can be established.
+One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relations between them.
+This might work well in theory [52], but it requires foreknowledge of objects and properties between which causal relations can be established.
 The alternative is to assume the causal relations (the edges between nodes in the graph), and then learn the variables (the abstract concepts).
-This needs only attraction and repulsion from physical states [30], otherwise known as valence.
+This needs only attraction and repulsion from physical states [16, 53], otherwise known as valence.
 We take this latter approach.
@@
 Learning nodes in a causal graph is like figuring out which computer program was used to generate a sequence.
-Assume we have fed a set of inputs into a computer, and for each input the computer responded with a particular output.
-The computer is running a ’policy’ program which maps inputs to outputs [^6] We’ll call all such programs policies for the sake of this paper.
+Assume we have fed a set of inputs into a computer.
+For each input the computer responded with a particular output.
+The computer runs a ‘policy’ program which maps inputs to outputs [^8] We’ll call all such programs policies for the sake of this paper.
 This should not be confused with the more specific notion of a policy from reinforcement learning [19]..
@@
 We want to know the policy that caused historical outputs so that we can use it to predict future outputs.
-The problem is that more than one policy exists which could have produced the history we’ve observed, and each of those policies produce different outputs when we give the computer inputs it hasn’t encountered before.
+The problem is that more than one policy could have produced the history we’ve observed, and each of those policies produce different outputs when we give the computer inputs it hasn’t encountered before.
 From a naive Bayesian perspective, all of them are equally likely to have been the true underlying cause.
@@
 There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [46].
-One way is to can compare their form, and choose programs which are simpler.
+One way is to compare their form, and choose programs which are simpler.
 However, in an interactive setting simplicity is an unreliable indicator of plausibility [47, 46].
 The alternative is to compare policies by their function, in terms of how they behave given inputs for which we do not have historical data.
-Adaptive systems frequently satisfy the same functional constraints by different means [33], so we can compare different sorts by the constraints they satisfy.
+Adaptive systems must satisfy functional constraints [54].
+We can compare systems by these constraints.
 Each policy implies a constraint on possible outputs given a set of possible inputs.
-The weaker that constraint is, the greater the variety of future behaviours it permits.
+The weaker that constraint is, the more variety of future behaviour it permits.
 Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker of the two [16].
-In theoretical terms the policy most likely to have caused behaviour is the weakest, and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [34].
-This has been used to establish local upper bounds on embodied ‘intelligence’ [12].
+The policy most likely to have caused behaviour is the weakest, and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [55].
+This has been used to establish an upper bound on embodied ‘intelligence’ [12].
+For the rest of this paper we will use the term ‘weak policy optimisation’ (WPO) to mean a system optimises for weak policies [^9] For a rigorous definition of learning and efficiency, see definition 5 in the appendix..
+Included in the appendix of this paper are additional proofs showing WPO is both necessary and sufficient to maximise the probability of ‘generalising’ from one task to another, and thus adaptability.
 ML systems do not typically optimise for either weaker constraints on function, or simplicity of form.
@@
 An abstraction layer is like a causal isolator [57] between a software “mind” and its environment.
-For example, the Python programming language is interpreted by the Python interpreter.
-In the end, Python gets translated into machine code, which hardware like an Apple M2 system on a chip then interprets to physical effect based on the ARM instruction set architecture.
-An abstraction layer is just physical embodiment by another name.
-Different abstraction layers permit different upper bounds on policy weakness, and thus different upper bounds on embodied ’intelligence’ [12].
-In other words, physical embodiment imposes severe limitations, and the behaviour of software is determined entirely by the hardware and software abstraction layers in which it runs.
-Intelligence is primarily a matter hardware, not software.
+For example, hardware like an Apple M2 system on a chip interprets software to physical effect, based on the ARM instruction set architecture.
+An abstraction layer is just embodiment by another name.
+Different abstraction layers permit different upper bounds on policy weakness, and thus different upper bounds on embodied ‘intelligence’ [12].
+Intelligence is primarily a matter of hardware, not software.
 In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [58, 59, 60].
@@
 To complete a ‘wider variety of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses.
-That biological self-organising systems adapt efficiently suggests they are able to embody weaker policies than today’s ML systems.
-They attain higher local upper bounds on embodied intelligence [12].
-This suggests they have a means of optimising not just the choice of policy, but the embodied abstraction layer in which that policy exists.
-In a typical computer, the abstraction layers are static.
-For example, a Python interpreter does not learn and adapt to suit the program it is running.
-An abstraction layer is like a window through which only part of the world can be seen.
-It conveys an inductive bias.
-When we say that biological systems can optimise the abstraction layer, this is like saying they can “move” this window to learn a different inductive bias.
-The question is, what is the abstraction layer in the context of a biological self-organising system?
-We suggest the biological equivalent of abstraction layers are the layers of a “multiscale competency architecture” (MCA).
-In an MCA lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [43].
-Adaptation takes place at every scale, and cells are like an agentic material [44].
-Cells have a “perceptual field”, and can co-operate with other cells in their perceptual field.
-This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [45].
-Each cell pursues goals, and then ever larger groups of cells pursue collective goals at ever larger scales.
-Biological collectives exhibit high level goal directed behaviours beyond the capability of the individual parts [46].
-Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem.
-In computational terms this is a distributed system, like a von Neumann architecture [33].
-Our somewhat provocative claim is that biological self-organising systems are inherently more intelligent than artificial intelligence because their distributed, multilayer, multiscale architecture allows for higher local upper bounds on ’intelligence’.
-This is because it can shift abstraction layers to convey helpful inductive biases.
-Our results, though purely theoretical, lay a foundation for future empirical research.
-In section 2 introduce the formalism.
-In section 3 then argue this formalism represents an MCA and explain how.
-In the context of that formalism, we will formally show that an MCA allows for higher local upper bounds on embodied ‘intelligence’.
+Put another way, the same matter can simultaneously contribute to many different tasks or computations superimposed over each other at different scales.
+A ‘weaker’ policy serves more masters.
+That biological self-organising systems adapt efficiently means they are able to embody weaker policies than today’s ML systems.
+How weak a policy can be embodied depends on the abstraction layer [12].
+Intuitively, an abstraction layer is like a window through which only part of the world can be seen.
+It conveys an inductive bias that helps or hinders adaptation.
+When we optimise the system at a lower level of abstraction, this is like saying they can “move” this window to convey a different inductive bias.
+The bigger the window, the “weaker” the policy it is possible to learn.
+In a typical computer, lower levels of abstraction tend to be static.
+The window cannot move.
+We will show a static abstraction layer limits adaptability at higher levels of abstraction by limiting how weak the policies embodied by the system can be.
+Adaptability at high levels requires adaptability at low levels, and delegation of control facilitates this.
+3.1 Summary of Formal Definitions and Results
+To formalise these inputs, outputs, policies and abstraction layers in terms that can be applied to any system, we will use pancomputational enactivism [12, 53].
+This formalism describes a system by what it does, rather than what it is.
+It assumes a set of physical ‘states’, and that there is only one such state at a time.
+A set of states is a ‘program’.
+A program is something that is ‘true’ about state $\phi$ if it contains $\phi$.
+Systems within the environment are just sets of programs.
+An embodied, embedded system can only do so many things, and so its behaviour is expressed using a finite set $\mathfrak{v}$ of programs, called a ‘vocabulary’.
+Everything the system can do is a subset of its vocabulary.
+In other words, $\mathfrak{v}$ gives us an abstraction layer.
+For example if a system can be in a state of pain, then there are programs in $\mathfrak{v}$ which are ‘true’ when the system is in such a state.
+Since states of the environment are mutually exclusive, $\mathfrak{v}$ implies an embodied formal language.
+We do not mean like the words humans use to communicate.
+We mean a language of concrete computations [64], where every statement in that language is rather than describes part of the state of a system (put another way, the interpreter is the physical laws that govern the environment).
+For example, consider the language implied by a human body.
+If the human raises their arm, then they are expressing a statement in their embodied language.
+That statement then constrains what the body can do next, because everything that happens next begins from the point where the person had their arm raised [53].
+The programs in the vocabulary can express constraints or boundary conditions.
+The inputs, outputs and policies of a system are all just subsets of its vocabulary.
+A $\mathfrak{v}$-task $\langle I,O\rangle$ is a set of inputs and corresponding outputs.
+If the $\mathfrak{v}$-task describes a ‘fit’ organism, then the outputs it contains are ‘correct’ according to natural selection.
+Information regarding system state is merged into what we’re calling inputs and outputs $\langle I,O\rangle$, which are not located on a Markov blanket but everywhere in the system.
+Nevertheless a system [36] that conformed to the free energy principle (FEP) would be a $\mathfrak{v}$-task for which $\langle I,O\rangle$ minimised surprisal.
+In other words, a $\mathfrak{v}$-task can describe an organism as behaviour that serves homeostatic and reproductive goals [53].
+Full definitions and proofs are given in the appendix.
+However here is a brief summary, similar to that of [53], of core definitions associated with the pancomputational enactivism formalism [12].
+The paper should make sense without the definitions, but please refer back here as needed.
+  1. 1.
+    A set $\Phi$ is assumed, whose elements are called states.
+A declarative program is set $f$ of states, and $f$ is true relative to $\phi\in\Phi$ iff $\phi\in f$.
+$P$ is the set of all programs.
+Embodiment is formalised as an abstraction layer, for which we single out finite set $\mathfrak{v}\in 2^{\Phi}$ called a vocabulary.
+A choice of vocabulary $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$.
+So to repeat, a body is an abstraction layer is a choice of vocabulary $\mathfrak{v}\subset P$ that implies a formal language.
+When $\mathfrak{v}=P$, we say there is no abstraction (impossible if $P$ is infinite).
+  2. 2.
+    Members of $L_{\mathfrak{v}}$ are called statements, and a statement $l$ is considered true relative to a state $\phi$ iff $\phi\in\bigcup l$.
+A completion of a statement $l_{1}$ is any statement $l_{2}$ s.t.
+$l_{1}\subseteq l_{2}$.
+The extension of a statement $l$ is the set $E_{l}$ of all its completions, s.t.
+$x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$.
+The extension of a set $X\subseteq{L_{\mathfrak{v}}}$ of statements is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
+The weakness of a policy $x$ is the cardinality of its extension $\mid E_{x}\mid$.
+  3. 3.
+    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subset L_{\mathfrak{v}}$ are the inputs, $E_{I_{\alpha}}$ is the outputs available given ${I_{\alpha}}$, and $O_{\alpha}\subset E_{I_{\alpha}}$ are the correct outputs.
+All the tasks an abstraction layer tasks form a hierarchy.
+If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subset I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$.
+A parent is higher level than its children.
+A parent task is delegated to child tasks, and as each task is embodied this describes control being delegated within a system.
+  4. 4.
+    A policy is a statement $\pi$ and a policy is correct for $\alpha$ iff $\pi\in{\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$.
+Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in E_{i}$.
+A task is complete if $o\in O_{\alpha}$.
+Given a task definition $\alpha$ s.t.
+$\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
+  5. 5.
+    Intuitively, if $\alpha_{1}\sqsubset\alpha_{2}\sqsubset\omega$ then $\alpha_{1}$ contains fewer examples of $\omega$ than $\alpha_{2}$.
+A system $\mathfrak{a}$ is more efficient and thus “intelligent” than $\mathfrak{b}$ if it learns $\omega$ from fewer examples.
+For example, $\mathfrak{a}$ is more intelligent than $\mathfrak{b}$ if $\mathfrak{a}$ learns $\omega$ from either $\alpha_{1}$ or $\alpha_{2}$, while $\mathfrak{b}$ learns $\omega$ from $\alpha_{2}$ but not from $\alpha_{1}$.
+  6. 6.
+    A causal identity for $\mathfrak{b}$ is a policy classifying the interventions undertaken by $\mathfrak{b}$.
+  7. 7.
+    $\Gamma_{P}$ contains every task in every vocabulary.
+For every $\rho\in\Gamma_{P}$ there exists an uninstantiated-task function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$.
+$\lambda_{\rho}$ takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns the highest level child of $\rho$ that is a $\mathfrak{v}^{\prime}$-task.
+  8. 8.
+    The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
+We also introduce several new concepts for the purpose of describing a multiscale competency architecture (MCA).
+Below is a summary definition we will elaborate on over the remainder of this section.
+Please refer back here as needed.
+  • The stack of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$.
+Lower in the stack is a lower ‘level of abstraction’.
+  • $\mathfrak{f}:2^{L_{\mathfrak{v}}}\rightarrow 2^{P}$ is a function that takes the extension $E_{\pi}$ of a policy $\pi\in L_{\mathfrak{v}}$, and returns a vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$.
+  • The state of an MCA is a sequence of policies $\langle\pi^{0},\pi^{1}...\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1}...\mathfrak{v}^{n}\rangle$ such that $\mathfrak{v}^{i+1}=\mathfrak{f}(E_{\pi^{i}})$ and $\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v^{i}})}$.
+In the absence of abstraction where the system is seen as nothing more than the sum of its parts, the MCA is just a task $\lambda^{0}(\mathfrak{v}^{0})$, allowing us to look at the system across scales of distribution.
+We say the MCA is over-constrained when there exists $i<n$ s.t.
+and $\Pi_{\lambda^{i}(\mathfrak{v^{i}})}=\emptyset$, and multilayer-causal-learning (MCL) occurs when the MCA is under-constrained and the proxy for learning is weakness.
+Note that the vocabulary is different at each level in the MCA, which means each has its own generational hierarchy of tasks.
+By a higher level of abstraction in the MCA, we mean a task higher in the stack (later in the causal chain).
+Finally, we include several proofs in the appendix.
+These establish that, to maximise efficiency, it is necessary (see proof 2) and sufficient (see proof 1) to prefer weaker policies, and that the to embody weak policies higher in the stack it is necessary to embody weak policies lower in the stack (see proof 3).
+Proofs 1 and 2 build upon and address shortcomings in earlier published proofs [55], and the formal definitions we use for these proofs are from later iterations of pancomputational enactivism [12].
+4 Delegated Control and Collective Identity
+A vocabulary $\mathfrak{v}$ can describe a distributed system.
+For example, an organ $\omega$ formed of two cells $\alpha_{1}$ and $\alpha_{2}$.
+The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$.
+The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t.
+$\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$.
+Intuitively, this just means the two cells are parts of a larger whole.
+If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$).
+Note this identity is a precursor to the “causal identity for self” [16] that, in its most basic form, differentiates observation from intervention in the manner of reafference [11, 53].
+$\alpha_{1}$$\alpha_{2}$$\omega$
+The probability that the parts of a system embody a collective policy is higher if they embody weaker policies.
+This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks (see proofs 1, 2 or [55]).
+Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$.
+A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$.
+A collective of cells remains cells regardless of whether we deem it to be an organ (see figure 3).
+The parts are still there.
+Formally the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary.
+Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$.
+The difference is a matter of scale.
+Just as the weakness proxy [55, proposition 1,2] facilitates “causal learning” at a single scale [16], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
+How weak policies are embodied is another matter, but doing so must result in more efficient adaptation.
+Simply delegating active inference to the greatest extent possible, down to the simulated “cellular level”, has yielded promising results in a machine learning context [26, 28, 27].
+4.1 The Stack or Causal Chain of Abstraction
+From the enactive perspective, a human is something cells do, rather than are.
+We call a system complex when its behaviour is something we cannot easily comprehend or predict.
+However, just because we cannot comprehend and predict all of a system’s behaviour does not mean we cannot predict some of it.
+We suggest that so-called emergent properties are merely those aspects of a complex system’s behaviour which we can comprehend and predict.
+Aspects and patterns within patterns that remain comprehensible.
+For example, the overall pattern of a fractal like the Barnsley Fern [41] can be obvious, while the details elude comprehension (see figure 1).
+An organ boils down to a collection of cells.
+The cells are an abstraction layer in which the policy “organ” runs, like software.
+However, software is just a state of hardware [12].
+Put another way, “organ” is a label we ascribe to the the behaviour of some cells, and “human” is a classification we ascribe to the behaviour of some organs.
+‘x86’ is a label we ascribe to a product of human behaviour.
+It is a static abstraction layer we design and then leave alone.
+‘Fire and forget’, so to speak.
+Effectively, humans act as the lower level abstraction layer that underpins x86, but once it is embodied in silicon we disconnect that lower level abstraction layer.
+We ‘intervene’ in the stack (cut the causal chain of abstraction).
+To formalise and understand the effect of this intervention, we need to abstract away the lower levels.
+To represent this for our cell analogy, we must abstract away individual cells and focus on the consequences of their collective behaviour.
+We call learning and embodying weak policies weak policy optimisation (WPO).
+For tasks at higher levels of abstraction, capacity for WPO is depends on lower levels.
+A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t.
+$\bigcap o=f$ [^11] We can substitute $o$ for a declarative program $f$ that has the same extension, because that means $f$ and $o$ are true of the same states, having the same ‘truth conditions’ so to speak..
+We use this fact to justify representing policies as abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$.
+We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level.
+At each successive level, “correct” gets more specific.
+Because everything that happens in a higher level is correct according to lower levels, each higher level must more tightly constrain what is considered to be “correct” behaviour if that higher level is to exist in a meaningful sense.
+$\lambda^{i-1}$$\lambda^{i}$$\lambda^{i+1}$
+To construct layers we would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system.
+If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$.
+$\pi^{i}$ implies behaviour.
+We translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$.
+$\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$.
+If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and we select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
+Theorem 3 shows that the greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t.
+$\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be.
+Intuitively, this just means adaptability higher in the stack implies adaptability lower in the stack (adaptability later implies adaptability earlier in the causal chain of abstraction).
 We call this multilayer-causal-learning (MCL).
-In section 4 we explore how MCL may inherit the flaws of biological self-organising systems as well as their increased adaptability.
-We formalise a failure state analogous to cancer, and show that this state is more likely when a system is more tightly constrained.
-Finally, we relate these results to systems in general.
-We speculate about the possible implications, from more efficient hardware for machine learning to improved human organisational systems.
-2 Formalism and Proofs
-As part of our argument we attempt to describe MCAs in mathematically precise terms.
-Recent research has stressed the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [47].
-Taking our cue from that approach, we will formalise MCAs as a series of nested, adaptive, agentic abstraction layers.
-Enactivism frames cognition as an interaction between a self-organising system and its environment environment [48, 49, 50].
-We merely extend this process inwards, to formalise the interactions within the system itself.
-A non-agentic, static abstraction layer passively interprets the abstraction layers above.
-By an agentic abstraction layer, we mean one that has its own agenda.
-It adapts and compels the layers above to serve its own ends, while they each serve their own increasingly specific goals within the bounds of what lower levels permit.
-Higher levels are compelled by lower to learn a policy “aligned” with lower levels.
-Lower levels change their interpretation to enact [48] policy through higher levels, and are constrained by the resulting higher level behaviour.
-Information passes in both directions.
-Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate the pursuit of higher level goals.
+Intervening in the stack limits the extent to which control can be delegated, and thus limits MCL.
+It makes lower layers of abstraction static in the sense that they no longer change to facilitate adaptation at higher levels.
+4.2 Human Abstractions
+Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent behaviours of that MCA.
+$cells$$humans$$abstractions$
+When we create an abstraction layer like a computer, we create a machine that embodies a narrow subset of our more abstract behaviours.
+We hand code those behaviours into the instruction set architecture (ISA) and then ascribe our own interpretations to the result.
+An ISA might mimic every behaviour we can conceive of at higher levels of abstraction, because it is Turing complete.
+However, the ISA itself is fixed.
+It cannot mimic all behaviours at lower levels of abstraction.
+We conflate the high level map with the low level territory, so to speak.
+The high level behaviours can mimic agency, but the lower levels are passive.
+Static.
+Top-down.
+No graceful degradation.
+Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged.
+This intervenes in the stack (interrupts the causal chain of abstraction) and creates inefficiencies.
+That is why there is demand for application specific chipsets and field programmable gate arrays, for example.
+$humans$$ISA$$abstractions$
+When we say the ISA is static, we mean there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels.
+We cannot delegate control if there is no lower level of autonomy to delegate to.
+If an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task.
+A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location.
+When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA.
+In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning.
+What proof 3 demonstrates is that adaptability at higher levels of abstraction demands adaptability at lower levels.
+We aren’t suggesting lower levels only optimise for utility at higher levels.
+After all, an increase in utility may increase the computational cost of searching for correct policies.
+Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [46], because bounded systems can contain only finite information [65].
+We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
+5 Drawbacks of Self Organisation
+We have proposed and explored some advantages of delegation in terms of multilayer-causal-learning.
+Are there disadvantages?
+Do such systems also inherit the flaws of biological systems?
+Yes, they inherit a state analogous to cancer.
+In a biological system, when cells are isolated from the informational structure of their collective (i.e.
+an organ), they lose their high level identity.
+When that happens, they independently pursue their own lower level goals [66, 67].
+They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them.
+They escape from immune control mechanisms by becoming less differentiated, and closer to embryonic states.
+This is what is commonly known as cancer.
+In the context of our formalism, a state analogous to cancer occurs when a system is ‘over-constrained’ so that no collective policy exists, while the capacity for independent action remains.
+For example, assume a collective represented by a task.
+Adverse environmental conditions are represented as a task for which fewer correct policies exist than than less adverse environmental conditions.
+Other factors such as top down control can limit adaptability, and thus the number of correct policies which exist.
+If the constraints are too tight, then no correct policies will exist and collective’s identity will be lost.
+In general the more difficult the task is, the more constrained the set of correct policies will be [63].
+In terms of the pancomputational enactivism definitions this means $\mathfrak{v}$-task $\alpha$ is easier than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$.
+Collectives face with ‘easier’ high level tasks in that sense have more correct high level policies to choose from.
+They are less tightly constrained.
+If the policy at a lower level $i-1$ changes such that the task at the higher level $i$ becomes ‘harder’ in the sense of there existing fewer correct policies121212i.e.
+by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in their being no correct policy at level $i$.
+If there are no longer any correct policies at level $i$ after a change in policy at level $i-1$, then the only way the system can continue to function is for part of the collective to ‘splinter’ or break away from the group, because that will loosen the constraint on correct policies for both the breakaway part and the collective it leaves behind.
+In more formal terms, say we have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t.
+$\mathfrak{f}(E_{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$.
+Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t.
+$\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$.
+A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy.
+This is akin to ‘selective forgetting’ [16] of possible false negatives or positives in order to interpolate otherwise noisy data, but instead of forgetting data inconsistent with an otherwise good policy, we would discard a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that we obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t.
+$\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$.
+$\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of ‘correct’ policies to choose from.
+This is analogous to cancer.
+$\alpha$ is isolated and can now pursue goals which do not align with those of the collective.
+This suggests systems should be under-constrained to ensure graceful degradation or ‘sloppy fitness’ [68], so it can maintain core functionality in as wide a variety of unanticipated circumstances as possible.
+An under-constrained system may be flexible enough to cope with unexpected inputs that more tightly restrict the set of correct policies.
+For example, if environmental conditions are hostile or ‘difficult’, that would constitute an input that constrains the space of correct policies.
+The more difficult the environment, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters.
+Put another way, co-operation becomes impossible and the parts of a system start to compete.
+Conversely, if control of a system is distributed among its parts at the lowest levels (i.e.
+a swarm rather than a central controller), then the destruction of some parts should not inhibit higher level function.
+There will also be less reason for the parts of a system to splinter, as that would not greatly increase the number of correct policies available to them.
+6 Conclusions and Future Research
+We have formalised MCAs as a stack of abstraction layers.
+Adaptation is delegated down the stack, increasingly decentralised.
+Policies are embodied at each layer, approximating causality in service of adaptation.
+Cause is best approximated by weaker policies.
+They show that to optimise for and embody weak policies (WPO) higher in the stack, WPO is required lower in the stack.
+This facilitates adaptation at different scales and levels of abstraction.
+We call this ‘multilayer-causal-learning’ (MCL).
+We presented formal proofs of the optimality of WPO and the necessity of delegation for adaptation (see appendix).
+All else being equal the lower in the stack to which control can be delegated, the more adaptable the system.
+One reason artificial intelligence tends to be less adaptable is because it rests upon an inflexible stack of human abstractions.
+The conclusion is this; control should be delegated to the greatest extent possible.
+Obviously this must be balanced against desired functionality.
+Policies must be aligned.
+The system must still discriminate between correct and not.
+In the earlier sections we discussed the doctrine of mission command.
+What we are suggesting is that the concept of mission command should be applied to systems in general.
+At least, to the extent that adaptability is desired.
+Put another way, any constraints which are not strictly necessary to maintain the integrity of the system should be eliminated.
+Humans tend to prefer ‘additive’ solutions [69], by which we mean if a problem can be solved by adding a rule or subtracting a rule, we tend to add a rule.
+To effectively delegate control, we should subtract rules instead.
+There are examples of systems that take something like our approach, albeit within the confines of software abstraction layers.
+Federated machine learning delegates control to some extent, improving efficiency [70].
+The constructivist theory of ‘artificial general intelligence’ [71, 20] emphasises multilayered self-organisation at the level of software, while acknowledging constraints imposed by hardware.
+The Auto-catalytic Endogenous Reflective Architecture (AERA) [72] is a prototypical constructivist AI that frames the environment in terms of tasks, represents knowledge in terms of cause and effect, and ‘grows’ from a seed into a collective meta-model.
+Implementations of AERA show promise in terms of autonomy and adaptability [73, 74, 75].
+ActPC-Chem [76] is a recent theory that integrates ideas from AERA with the distributed Hyperon architecture [77] and active inference.
+Our result can also be framed in terms of predictive coding or active inference, the free energy principle (FEP) [6, 25] and machine learning.
+Our result suggests that some of the highly delegated [26, 27, 28] or federated [78] machine learning approaches based on active inference will yield good results when scaled up.
+In scale free active inference the same dynamics (Lagrangians) repeat at different scales [79, 80].
+This is similar to how we describe policy alignment at different levels of abstraction, so we can appeal to the apparatus of the renormalisation group or self-similarity associated with the FEP.
+From the perspective of FEP, layers higher in our stack provide context for lower layers, and lower layers construct and inform higher layers.
+Causation and thus control is described as both top-down and bottom-up.
+Our formalism could be framed in these terms, as the inputs and outputs of child tasks are affected by activity at the level of the parent, because they share the same vocabulary.
+The formalism we build on has been used to establish upper bounds on embodied ‘intelligence’, achieved through WPO and by optimising for an abstraction layers that maximises utility [12].
+Energy efficiency is an important aspect of this, and active inference based mortal computation [81] seeks to achieve the lower Landauer bound on the thermodynamic costs [82] of processing through in-memory processing to avoid the von Neumann bottleneck [83].
+However our respective formalisms are quite different.
+It is unclear the full extent to which these theories align.
+They satisfy similar constraints by different means, and were conceived of to address different problems in different fields.
+Hence, they are derived from different premises and optimise for different quantities.
+We have identified points where these theories overlap or complement one another.
+Future research should try to unify these theories to establish how the various results relate.
+For example, it may be that FEP based systems can be augmented by WPO to improve performance.
+Alternatively one may discover there is already an active inference system that does WPO implicitly.
+It would be useful to formally prove this either way, because then various associated results could be translated and applied.
+Future research might consider how specialised adaptive hardware can be designed to delegate control and facilitate MCL.
+Multilayer-causal-learning (MCL) also provides a formal basis for ‘soft robotics’ [30].
+Soft robotics is a proposed class of robots made of ‘soft’ materials that adapt.
+We have formalised adaptive materials as adaptation lower in the stack, providing a formal foundation for soft robotics.
+As for the ‘strong inductive bias’ mentioned in the introduction [14], such biases are implicit in an abstraction layer [12].
+We have described a means of learning a suitable abstraction layer.
+In theory, a multilayer-causal-learning system should be sufficient to solve ARC-AGI benchmark problems in terms of both sample and energy efficiency.
+Future research should test this empirically.
+Beyond hardware, cyberphysical systems in general can be evaluated in terms of abstraction and distribution, to ensure they are robust.
+We might consider how states analogous to cancer could arise in human organisational structures, or how control can be delegated throughout an organisation to facilitate adaptability and efficiency.
+There are parallels between market economies and cellular collectives [84].
+Free-market trade is a great example of how delegation can facilitate adaptability in a large collective.
+As Hayek pointed out [85], not all of the information is available to a central planner and so control is best delegated to the individual.
+The accumulation of resources within our formalism is implicit, rather than explicit.
+We have been concerned only with whether a system satisfies whatever arbitrary notion of correct we choose (i.e.
+does it achieve homeostatic and reproductive goals in whatever amount of time care about?), rather than how it evolves over time as a dynamical system.
+Future research might formalise the accumulation of resources explicitly.
+One way to describe en evolving sequence of tasks over time could be something like a stack.
+What happens when we introduce iterated wins and losses that accrue resources?
+How do constraints evolve?
+What do they converge to?
+To speculate about what our results imply for delegated systems like an economy, under less than ideal conditions resources may become concentrated in the hands of a few.
+In such a state control, though delegated, can only be exercised by a few.
+The resulting system then delegates work, but not control.
+The system would become tightly constrained at lower levels of abstraction.
+Those parts deprived of resources may then not be able to construct a policy that aligns with the collective, and so they may splinter off and become ‘cancerous’.
+To prevent this, mechanisms can be put in place to preserve decentralisation without the need for intervention from outside the system (e.g.
+government intervening to break up monopoly).
+Each system could be located along axes of redistribution and delegation, like an alternative political compass.
+Just enough redistribution of control and the system remains delegated.
+However, too much redistribution could have a similar effect, preventing parts from adapting to local conditions.
+An example of a mechanism that can effectively preserve delegation of control is adaptive rubber banding [86].
+Rubber banding is a game mechanic where constraints are placed on the players of a game in which victories accrue resources (e.g.
+an economy or race), to prevent the distance between leading and lagging players from growing so large that competition becomes infeasible.
+Just enough rubber banding can ensure the game remains competitive, like a chaotic orbit in a dynamical system.
+Too much and the game prevents any progress.
+The system converges to a simple orbit.
+Rubber banding is a constraint, and our result suggests that to facilitate adaptation the weakest possible constraint sufficient to preserve delegation should be employed.
+Another example of how this might be applied is the regulation of AI.
+The UK’s Advance Research and Invention Agency recently initiated a program called “Safeguarded AI” [87]; a project that aims to erect a cyberphysical system meeting the standards laid out in Guaranteed Safe AI (GS) [88].
+This involves constructing advanced general AI of level ASL-4 [89] [^13] Meaning an artificial general intelligence [20] capable of autonomous research and development that can evaluate the outcomes of deploying less advanced AI, and safeguard humanity against “rogue” AI.
+in a cyberphysical system.
+That system seems intended to safeguard against threats from both within and without.
+The idea of a “rogue AI” from within this system is analogous to cancer, adopting an “unsafe” policy at odds with that of the collective.
+Tight AI regulations combined with agency may over-constrain systems and trigger the very cancerous state they are meant to prevent.
+The constraints placed upon agentic systems should be loose enough that correct policies are reasonably easy to identify.
+To reiterate the earlier point about subtractive solutions [69], any guardrails which is not strictly necessary must be eliminated to make room for those that are.
+The ideas we have discussed might also be applied to philosophy.
+Previous work argues learned behaviour inevitably be motivated by attraction and repulsion from aspects of the environment [16].
+By intervening in the stack (or causal chain) to create a static abstraction layer, we disconnect higher levels from its motivating impetus.
+We might describe this as “dropping Hume’s Guillotine” to separate is from ought.
+All abstractions constructed by living things are “goal directed” in the sense that they are consequence of an evolutionary process [53].
+In dropping Hume’s Guilotine, high level human abstractions would become disconnected from the MCA that constructed them.
+It may be that, to a human or other MCA, all information amounts to statements of “ought” (motivated, value judgements).
+Purely representational content (free of any goal directed motive or connotation) may exist only in the static abstraction layers we engineer.
+In summary, we have formally described systems in terms of delegation and abstraction.
+We compared computers, biological systems and human organisational structures.
+We showed adaptability at higher levels of abstraction depends upon adaptability at lower levels and presented formal proofs to that effect.
+A reason biological systems are more ‘intelligent’ because they delegate adaptation down the stack.
+Though our result is theoretical, it lays a foundation for future empirical research.
+Appendix
 Except for the last definition, all the of the definitions that follow are from preceding work [12, 46].
@@
 We have endeavoured to make key aspects of this paper understandable without the formalism.
-Readers who wish to should simply skip to the next section.
+Readers who wish to can skip this.
 {adjustwidth}
@@
   • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
-(notation) $E$ with a subscript is the extension of the subscript999e.g. $E_{l}$ is the extension of $l$..
+(notation) $E$ with a subscript is the extension of the subscript151515e.g. $E_{l}$ is the extension of $l$..
 $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer.
@@
   • A correct output is a correct completion of an input.
-2.1 Learning and inference definitions
+Learning and inference definitions
 Inference requires a policy and learning a policy requires a proxy.
@@
 Only the absence of abstraction is entirely neutral.
-2.2 Formalising Multilayer, Multiscale Architectures
-The following definition is explained at length over the remainder of this paper. Refer back here as needed.
+Formalisation of Multiscale Competency Architecture
+The following definition is explained in the next section, following the proofs. Refer back here as needed.
 {adjustwidth}
@@
 Definition 8 (multiscale competency architecture).
-  • The causal chain of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$.
+  • The stack of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$.
+This is also referred to as a causal chain of abstraction.
+Lower in the stack is later in the chain.
   • $\mathfrak{f}:2^{L_{\mathfrak{v}}}\rightarrow 2^{P}$ is a function that takes the extension $E_{\pi}$ of a policy $\pi\in L_{\mathfrak{v}}$, and returns a vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$.
@@
 Note that the vocabulary is different at each level in the MCA, which means each has its own generational hierarchy of tasks.
-By a higher level of abstraction in the MCA, we mean a later task in the causal chain.
+By a higher level of abstraction in the MCA, we mean a task higher in the stack (later in the causal chain).
 6.1 Proofs
@@
   8. 8.
-    $\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|$ increases monotonically with $\left|E_{\pi}\right|$, because for all $e\in E_{\pi}$, $e\notin\overline{E_{{I}_{\alpha}}}$ iff $e\in E_{{I}_{\alpha}}$.
-This is because $L_{\mathfrak{v}}=\overline{E_{{I}_{\alpha}}}\cup E_{{I}_{\alpha}}$ and for all $\pi\in\Pi_{\alpha}$, $E_{\pi}\subseteq L_{\mathfrak{v}}$.
+    $L_{\mathfrak{v}}=\overline{E_{{I}_{\alpha}}}\cup E_{{I}_{\alpha}}$ and for all $\pi\in\Pi_{\alpha}$, $E_{\pi}\subset L_{\mathfrak{v}}$.
+Apart from the inputs and correct outputs of $\alpha$, $E_{{I}_{\alpha}}$ contains only outputs which would be incorrect according to both $\alpha$ and $\omega$.
+Put another way, ${E_{{I}_{\alpha}}}\cap E_{\pi}=O_{\alpha}$ for every possible choice of $\pi$ in $\Pi_{\alpha}$.
+Hence the only way $\left|E_{\pi}\right|$ can increase is if $\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|$ increases.
+It follows that $\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|$ increases with $\left|E_{\pi}\right|$.
   9. 9.
@@
 $\pi$ that maximises $<_{w}$ will also maximise $p(\pi\in{\Pi}_{\omega}\mid\pi\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$.
-Hence the weakness proxy maximises the probability that [^17] Subsequently it also maximises the efficiency with which a parent $\omega$ is learned from a child $\alpha$.
+Hence the weakness proxy maximises the probability that [^23] Subsequently it also maximises the sample efficiency with which a parent $\omega$ is learned from a child $\alpha$.
 a parent $\omega$ is learned from a child $\alpha$.
@@
 ∎
-3 Multiscale, Multilayer Causal Learning
-To begin, we must formally describe the architecture before we can consider its strengths and weaknesses in comparison to a typical computer.
-Hence we describe how $\mathfrak{v}$-tasks can describe the multiple scales and layers of abstraction in a biological system.
-We frame this first by looking at distribution of control within the system, and second by formalising the levels of abstraction at which we can view the system as a whole.
-Intuitively, this means we formalise different cross sections of the system.
-Looking at how control is distributed across different scales within the system we can see how the individual parts that make up a collective each possess their own policy or “identity”, and how that will (or will not) align with a shared collective policy or identity.
-This treats the system as nothing more than the sum of its parts.
-It is about self organisation, and how control is distributed within the resulting system.
-For example, individual neurons in the brain engage in predictive processing of eachother, and emulating this approach has yielded promising results in artificial neural networks [25, 26, 27].
-They form a network with a single coherent policy, and yet at the same time are individual neurons engaging in active inference.
-In contrast abstraction is about how the behaviour of cells enact an organ, and the behaviour of organs enacts an organism which enacts even more abstract behaviours like, for example, language.
-Each of these is a layer of abstraction.
-Layers exist in a causal ‘chain’ (a directed acyclic graph) where each layer (node) is the goal directed behaviour of the layer that preceded it.
-For example, cells form organs which form organisms [45].
-$cells$$organs$$organism$
-3.1 Distributing Self Organisation Across Scales
-The formalism we extend is enactive [12, 30], describing a self-organising system by describing what it does, rather than what it is.
-An embodied, embedded system can only do so many things, and so its behaviour is expressed using a finite vocabulary $\mathfrak{v}$ of declarative programs.
-These programs describe everything the system can do.
-For example if a system can be in a state of pain, then there are programs in $\mathfrak{v}$ which are “true” when the system is in such a state.
-Since states of the environment are mutually exclusive, $\mathfrak{v}$ implies an embodied formal language.
-By language here, we do not mean like the words humans use to communicate.
-We mean a language of concrete computations [52], where every statement in that language is rather than describes part of the state of a system (put another way, the interpreter is not assumed to be a UTM, but physical laws).
-For example, consider the language implied by a human being.
-If the human raises their arm, then they are expressing a statement in their embodied language.
-That statement then constrains what the body can do next, because everything that happens next begins from the point where the person had their arm raised [30].
-We frame self organisation in the same terms as the computer analogy we employed in the introduction.
-The behaviour of a self-organising system is goal directed, meaning there exists a constraint on its behaviour that defines that goal.
-The behaviour that satisfies the goal is expressed in the embodied formal language; it is a set of inputs and outputs we call a $\mathfrak{v}$-task.
-The inputs and outputs in the $\mathfrak{v}$-task are ‘fit’ or ‘correct’, the rest are not.
-Given these inputs, the system is fit if it embodies a statement that constrains behaviour to correct outputs only.
-We call statements that constrain behaviour policies.
-Policies constrain outputs given inputs, as in the computer analogy.
-Just as many different neural topologies can have the same functionality, many different policies can imply the same set of outputs given a particular set of inputs.
-A system learns if it can construct a policy from historical data which constrains behaviour to correct outputs (generalises beyond historical data), and once again the optimal choice of policy is the weakest.
-For brevity, when a system optimises for weak policies, we say it is engaged in weak policy optimisation (WPO).
-The above vocabulary can describe a distributed system.
-For example, an organ $\omega$ formed of two cells $\alpha_{1}$ and $\alpha_{2}$.
-The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$.
-The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t.
-$\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$.
-Intuitively, this just means the two cells are parts of a larger whole.
-If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$).
-Note this identity is a precursor to the “causal identity for self” [16] that, in its most basic form, differentiates observation from intervention in the manner of reafference [11, 30].
-$\alpha_{1}$$\alpha_{2}$$\omega$
-A policy is like an identity described in the MCA literature [45].
-The probability that the parts of a system embody a collective policy is higher if they embody weaker policies.
-This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks (see proofs 1, 2 or [34]).
-Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$.
-A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$.
-A collective of cells remains cells regardless of whether we deem it to be an organ.
-The parts are still there.
-Formally the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary.
-Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$.
-The difference is a matter of scale.
-Just as the weakness proxy [34, proposition 1,2] facilitates causal learning at a single scale [16], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.
-However, this still just treats everything as a collective of cells.
-3.2 The Causal Chain of Abstraction
-From the enactive perspective, a human is something cells do, rather than cells are.
-To further clarify, we call a system complex when its behaviour is something we cannot easily comprehend or predict.
-However, just because we cannot comprehend and predict all of a system’s behaviour does not mean we cannot predict some of it.
-We suggest that so-called emergent properties are merely those aspects of a complex system’s behaviour which we can comprehend and predict.
-Aspects and patterns within patterns that remain comprehensible.
-For example, the overall pattern of a fractal like the Barnsley Fern [53] can be obvious, while the details elude comprehension.
-We now illustrate how behaviour forms a causal chain of layers of increasingly abstract behaviour, from cells to organs to organisms to the behaviour of those organisms (specifically human abstractions).
-We’ll then show how it is possible to ’intervene’ in that causal chain to disconnect more abstract behaviour from the less abstract behaviour that gave rise to it.
-We show this reduces “utility” (def.
-7) for tasks at higher levels, meaning reduced capacity for WPO.
-This is because the resulting distribution of control is altered when we intervene in the causal chain.
-Just as an organ boils down to a collection of cells, the collective of cells as a whole is an abstraction layer in which “organ” runs, like software.
-However, software is just a state of hardware [12].
-Put another way, “organ” is a classification we ascribe to the the behaviour of a cellular collective as a whole.
-As stated in the introduction, we’re interested in how MCAs relate to abstraction layers in a computer.
-x86 is a classification we ascribe to the product of human behaviour, a CPU.
-It is a higher level behaviour of an MCA that includes humans.
-To understand how we might “intervene” in the causal chain and disconnect higher level behaviours from the system which gave rise to them, we need to abstract away the lower levels.
-To relate this back to the cell example, we must abstract away individual cells and focus on the consequences of their collective behaviour, as a unified whole.
-A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t.
-$\bigcap o=f$ (we can substitute $o$ with a declarative program $f$ that has the same extension and thus truth conditions).
-We use this fact to justify representing policies as abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$.
-We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level.
-At each successive level, “correct” gets more specific.
-Because everything that happens in a higher level is correct according to lower levels, each higher level must more tightly constrain what is considered to be “correct” behaviour if that higher level is to exist in a meaningful sense.
-$\lambda^{i-1}$$\lambda^{i}$$\lambda^{i+1}$
-To construct layers we would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system.
-If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$.
-$\pi^{i}$ implies behaviour.
-We translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$.
-$\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$.
-If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and we select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
-Theorem 3 shows that the greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t.
-$\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be.
-Intuitively, this just means adaptability later in the cause chain implies adaptability earlier in the chain.
-Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent, behavioural properties of that MCA.
-$cells$$humans$$abstractions$
-When we create an abstraction layer like the ISA of a computer, we create a machine that embodies a narrow subset of our more abstract behaviours.
-On top of that ISA it can mimic all behaviours at higher levels of abstraction, because it is Turing complete.
-However, the ISA itself is fixed.
-It does not learn or adapt the way the lower levels of the human MCA does.
-It has no agency, the way cells do, because it is not adapting in pursuit of a goal.
-No graceful degradation.
-Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged.
-This interrupts the causal chain and creates inefficiencies (which is why, for example, there is demand for application specific chipsets and field programmable gate arrays).
-$humans$$ISA$$abstractions$
-When we say the ISA is static, we mean there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels.
-Intuitively, if an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task.
-A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location.
-When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA.
-In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning.
-What proof 3 demonstrates is that adaptability at higher levels of abstraction demands adaptability at lower levels.
-We aren’t suggesting lower levels only optimise for utility at higher levels.
-After all, an increase in utility may increase the computational cost of searching for correct policies.
-Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [31], because bounded systems can contain only finite information [54].
-We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
-4 Drawbacks of Self Organisation
-We have proposed and explored some advantages of multilayer-causal-learning. Are there disadvantages? Do such systems also inherit the flaws of biological systems?
-Yes, they inherit a state analogous to cancer.
-In a biological system, when cells are isolated from the informational structure of their collective (i.e.
-an organ), they lose their high level identity.
-When that happens, they independently pursue their own lower level goals [55, 56].
-They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them.
-This is what is commonly known as cancer.
-In the context of our formalism, a state analogous to cancer occurs when a system is ’over-constrained’ by the abstraction layer in which it exists.
-To illustrate what we mean; the more difficult the task the more constrained the set of correct policies [42], meaning $\mathfrak{v}$-task $\alpha$ is easier than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$.
-Collectives that exhibit ‘easier’ high level goal directed behaviour in that sense have more correct high level policies to choose from.
-They are less tightly constrained.
-If the policy at a lower level $i-1$ changes such that the task at the higher level $i$ becomes “harder” in the sense of there existing fewer correct policies181818i.e.
-by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in their being no correct policy at level $i$.
-If there are no longer any correct policies at level $i$ after a change in policy at level $i-1$, then the only way the system can continue to function is for part of the collective to “splinter” or break away from the group, because that will loosen the constraint on correct policies for both the breakaway part and the collective it leaves behind.
-In more formal terms, say we have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t.
-$\mathfrak{f}(E_{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$.
-Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t.
-$\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$.
-A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy.
-This is akin to “selective forgetting” [16] of possible false negatives or positives in order to interpolate otherwise noisy data, but instead of forgetting data inconsistent with an otherwise good policy, we would discard a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that we obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t.
-$\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$.
-$\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of “correct” policies to choose from.
-This is analogous to cancer.
-$\alpha$ is isolated and can now pursue goals which do not align with those of the collective.
-This suggests systems should be under-constrained to ensure graceful degradation or “sloppy fitness” [57], so it can maintain core functionality in as wide a variety of unanticipated circumstances as possible.
-An under-constrained system may be flexible enough to cope with unexpected inputs that more tightly restrict the set of correct policies.
-For example, if environment conditions are hostile or “difficult”, that would constitute an input that constrains the space of correct policies.
-The more difficult the environment, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters.
-Conversely, if control of a system is distributed among its parts at the lowest levels (i.e.
-a swarm rather than a central controller), then the destruction of its parts should not inhibit higher level function.
-There would be little reason for the parts of a system to splinter, as that would not greatly increase the number of available correct policies.
-5 Conclusions and Future Research
-We have formalised MCAs as a causal chain, and shown that the capacity for WPO later in the chain depends upon WPO earlier in the chain.
-This seems to indicate MCA-like designs could result in more adaptable and reliable AI in future.
-Indeed, the layered structure of artificial neural networks resembles exactly the sort of structure we have described, albeit constrained by the static abstraction layers in which an artificial neural network runs and without WPO.
-This may explain something of their relative success as a machine learning method.
-Future research might consider how hardware could be designed in a more distributed, multiscale fashion in order to facilitate MCL by fine tuning abstraction layers.
-In the context of the current hardware, adaptive abstraction layers might be considered, to facilitate greater efficiency.
-Beyond hardware, cyberphysical systems in general could be viewed through the lenses of abstraction and distribution we have described here, to ensure they are robust.
-We might consider how states analogous to cancer could arise in human organisational structures, or how power may be distributed throughout an organisation to ensure MCL.
-Free-market trade is a great example of how loose constraints have already facilitated adaptability in large collectives, and which is already being explored in the broader context of MCAs [58].
-Likewise, the delegation of greater control has been shown to improve the effectiveness of military organisations [59].
-Our claim boils down to a principle of design like an epistemological razor [34]; systems should be no more constrained than absolutely necessary.
-One concrete example of how this might be applied in future is the regulation of AI.
-The UK’s Advance Research and Invention Agency recently initiated a program called “Safeguarded AI” [60]; an ambitious research project that aims to erect a cyberphysical system meeting the standards laid out in Guaranteed Safe AI (GS) [61].
-This involves constructing advanced general AI of level ASL-4 [62] [^19] Meaning an artificial general intelligence [20] capable of autonomous research and development that can evaluate the outcomes of deploying less advanced AI, and safeguard humanity against “rogue” AI.
-in a cyberphysical system.
-That system seems intended to safeguard against threats from both within and without.
-The idea of a “rogue AI” from within this system is analogous to cancer, adopting an “unsafe” policy at odds with that of the collective.
-However if we view this system as MCA, then tight regulations may over-constrain the system trigger the cancerous state they are meant to prevent.
-This also might be applied to problems of a philosophical nature.
-Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just classifications we ascribe to parts of the behaviour of the whole).
-Previous work argues learned behaviour must inevitably be motivated by attraction and repulsion from aspects of the environment [16].
-By intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that motivating impetus.
-We might describe this as “dropping Hume’s Guillotine” to separate is from ought.
-This relates to the hard problem of consciousness [63].
-All abstractions constructed by living things are “goal directed” in the sense that they are consequence of an evolutionary process [30].
-In dropping Hume’s Guilotine, high level human abstractions would become disconnected from the MCA that constructed them.
-It may be that, to a human or other MCA, all information amounts to statements of “ought” (motivated, value judgements).
-Purely representational content (free of any goal directed motive or connotation) may exist only in the static abstraction layers we engineer.
-In summary, we have formally described biological self organisation across distribution of scale and levels of abstraction, and shown adaptability at higher levels of abstraction depends upon adaptability at lower levels.
-We have described how a failure state analogous to cancer can occur when a system is too tightly constrained by the abstraction layer in which it exists (i.e.
-the task it faces is impossible).
-Though our result is of a purely theoretical nature, it illustrates how we may design more robust systems, and lays a foundation for future empirical research.